### v2

# The Future Outcome Reasoning and Confidence Assessment Benchmark

Abstract
Forecasting is an important task in many domains, such as technology and economics. However existing forecasting benchmarks largely lack comprehensive confidence assessment, focus on limited question types, and often consist of artificial questions that do not align with real-world human forecasting needs. To address these gaps, we introduce FOReCAst (Future Outcome Reasoning and Confidence Assessment), a benchmark that evaluates models’ ability to make predictions and their confidence in them. FOReCAst spans diverse forecasting scenarios involving Boolean questions, timeframe prediction, and quantity estimation, enabling a comprehensive evaluation of both prediction accuracy and confidence calibration for real-world applications.
The Future Outcome Reasoning and Confidence Assessment Benchmark
Zhangdie Yuan University of Cambridge zy317@cam.ac.uk                        Zifeng Ding University of Cambridge zd320@cam.ac.uk                        Andreas Vlachos University of Cambridge av308@cam.ac.uk
Type Question Resolution Confidence
Boolean Question Will a Frontier AI lab be established in China by 2026? Yes 0.73
Timeframe Prediction When will OpenAI announce GPT-5? 2024-08-01 0.85
Quantity Estimation How many spacecrafts will land on the moon in 2025? 3 0.65
Table 1: Examples of forecasting questions with their resolutions and confidence scores.
1 Introduction
Recent advances in large language models (LLMs) have significantly improved their performance across a wide range of natural language processing (NLP) tasks. Alongside these developments, various benchmarks and datasets have been introduced to effectively assess the capabilities of LLMs, particularly in terms of knowledge and reasoning (Zellers et al., 2019; Guo et al., 2023). Fact-based benchmarks, like TruthfulQA (Lin et al., 2022) evaluate LLMs based on factual correctness, focusing on tasks like retrieving and verifying facts that are known.
Forecasting is a crucial yet challenging task across various domains, including technology, economics, and public policy. Unlike tasks that rely on retrieving and verifying existing knowledge, forecasting requires predicting plausible outcomes for future events, often under uncertainty and incomplete information. This makes forecasting particularly difficult, as models must infer trends, assess probabilities, and adapt to new information. Several datasets have been introduced to evaluate LLMs’ forecasting capabilities. ForecastQA (Jin et al., 2020) uses a multiple-choice format where models predict future outcomes, but it lacks confidence assessment. AutoCast (Zou et al., 2022) incorporates confidence intervals, however its confidence estimates are not designed for forecasting. Other datasets such as ExpTime (Yuan et al., 2024a) are artificially generated from structured data, focusing on explainable event forecasting based on temporal knowledge graphs.
All of these aforementioned benchmarks ignore a crucial aspect of forecasting: confidence evaluation. Confidence plays a central role in forecasting, as predictions about unresolved events inherently lack definitive correctness at the time of evaluation. Predictions made with absolute certainty are undesirable, even if they ultimately prove to be correct, because they fail to account for the uncertain nature of future events. Moreover, miscalibrated confidence can lead to poor decision-making: overconfident yet incorrect forecasts may result in costly errors, while underconfident but accurate predictions can erode trust in the model. Therefore, well-calibrated confidence scores are as crucial as the accuracy of the predictions themselves.
To address these gaps, we present FOReCAst: Future Outcome Reasoning and Confidence Assessment. FOReCAst focuses on three distinct types of forecasting questions, shown in Table 1: (1) Boolean questions, such as "Will there be a Frontier AI lab in China before 2026?"; (2) Timeframe Prediction, such as "When will OpenAI announce GPT-5?"; and (3) Quantity Estimation, such as "How many spacecrafts will land on the moon in each of the following years?" We conduct experiments using a range of models differing in size, training objectives, and cutoff times, and explore multiple methods for estimating model confidence. Our results reveal that forecasting remains highly challenging for current LLMs, particularly in confidence evaluation, with no direct correlation between prediction performance and confidence calibration, and while larger models sometimes improve performance, the effect is inconsistent.
2 FOReCAst: Problem Formulation
System responses in FOReCAst consist of (1) a prediction answering a question given the available information and (2) a confidence score in the prediction. This
ensures a comprehensive assessment of forecasting, accounting for both correctness and confidence calibration.
Questions belong to three types. (1) Boolean Questions, which ask yes/no questions about the occurrence of future events (sometimes within a certain timeframe). Boolean questions are simple to evaluate, and they can still be surprisingly challenging  (Clark et al., 2019). (2) Timeframe Prediction, which requires predicting a specific timeframe for an event, and are essential for applications where knowing whether an event will happen or not without a timeframe is insufficient. (3) Quantity Estimation, which involves providing numerical estimates related to future events, e.g., economic indicators or trends.
Formally, let $Q$ represent a question about a future event, and let $M$ denote a system with access to information up to time $t_{train}$ (e.g., the system’s knowledge cutoff point). The objective is for $M$ to produce an answer $A$ in $\mathcal{A}$ and an associated confidence score $C$, where:
$M(Q)\to(A,C),\quad C\in[0,1].$ (1)
$A=\arg\max_{a\in\mathcal{A}}P(X=a|Q,\mathcal{K}(t_{train})).$ (2)
Here, $\mathcal{K}(t_{train})$ represents the knowledge accessible to the model up to time $t_{train}$. The answer space $\mathcal{A}$ depends on the type of forecasting question: for Boolean Questions, $\mathcal{A}=\{\text{Yes},\text{No}\}$; for Timeframe Prediction, $\mathcal{A}$ consists of a single date in the YYYY-MM-DD format; and for Quantity Estimation, $\mathcal{A}=\mathbb{R}$, representing real numbers.
3 Evaluating Predictions and Confidence
Evaluating Boolean Questions
For Boolean questions, where the answer space is $\mathcal{A}=\{\text{Yes},\text{No}\}$, prediction performance is evaluated using standard classification metrics, including accuracy and F1-score.
Confidence calibration is assessed using a modified version of the Brier score (Brier, 1950), which measures the mean squared error between predicted confidence and the gold confidence provided in the dataset, which we assume is provided, and
represents the likelihood of an event occurring. The modified Brier score is defined as:
$\text{Brier}=\frac{1}{N}\sum_{i=1}^{N}(C^{\text{pred}}_{i}-C^{\text{gold}}_{i})^{2},$ (3)
where $C^{\text{pred}}_{i}$ is the model’s predicted confidence and $C^{\text{gold}}_{i}$ is the gold confidence. This modification ensures that models are evaluated based on their ability to match the likelihood of an event. Lower Brier scores indicate better calibration, reflecting how well the predicted confidence aligns with the likelihood of the event.
Evaluating Timeframe Prediction
For timeframe prediction, where the answer space consists of specific dates in the YYYY-MM-DD format, predictive accuracy is measured using absolute day error (ADE). Given a predicted date $D^{\text{pred}}_{i}$ and the gold date $D^{\text{gold}}_{i}$, we compute the normalized error as:
$E_{i}^{\text{ADE}}=\frac{2}{1+e^{-\alpha|D^{\text{pred}}_{i}-D^{\text{gold}}_{i}|}}-1,$ (4)
where $\alpha$ is a scaling factor that controls how sharply large errors are penalized. This transformation ensures that extreme deviations do not disproportionately dominate the evaluation.
For confidence calibration, rely on the Continuous Ranked Probability Score (CRPS) (Matheson and Winkler, 1976), which is a generalisation of the Mean Absolute Error to probabilistic forecasts, and extend it to compare the predicted probability distribution with a gold distribution. Specifically, we assume that both the predicted and the gold confidence predictions follow Gaussian distributions, namely $\mathcal{N}(D^{\text{pred}}_{i},\sigma^{\text{pred}}_{i})$ and $\mathcal{N}(D^{\text{gold}}_{i},\sigma^{\text{gold}}_{i})$ respectively, where the standard deviations are computed as:
$\sigma^{\text{pred}}_{i}=\sigma_{\max}\cdot(1-C^{\text{pred}}_{i})+\sigma_{\min}\cdot C^{\text{pred}}_{i},$ (5)
$\sigma^{\text{gold}}_{i}=\sigma_{\max}\cdot(1-C^{\text{gold}}_{i})+\sigma_{\min}\cdot C^{\text{gold}}_{i}.$ (6)
Here, $C^{\text{pred}}_{i}$ is the model’s predicted confidence for the $i$th question, and $C^{\text{gold}}_{i}$ is the corresponding gold confidence provided in our dataset. The parameters $\sigma_{\max}$ and $\sigma_{\min}$ define the upper and lower bounds for the standard deviation. Intuitively, when confidence is low ($C\approx 0$), uncertainty is high, leading to $\sigma\approx\sigma_{\max}$, while when confidence is high ($C\approx 1$), uncertainty is low, resulting in $\sigma\approx\sigma_{\min}$. We then compute the CRPS as the integrated squared difference between the cumulative distribution functions (CDFs) of the predicted and gold distributions:
$\text{CRPS}=\frac{1}{N}\sum_{i=1}^{N}\int\left(F^{\text{pred}}_{i}(d)-F^{\text{gold}}_{i}(d)\right)^{2}\,\mathrm{d}d,$ (7)
where $F^{\text{pred}}_{i}$ and $F^{\text{gold}}_{i}$ denote the CDFs of the predicted and gold Gaussian distributions, respectively. A lower CRPS indicates better calibration, as it reflects a closer match between the predicted uncertainty and the uncertainty as specified by the gold confidence.
Evaluating Quantity Estimation
For quantity estimation, where the answer space consists of non-negative real numbers ($\mathcal{A}=\mathbb{R}_{\geq 0}$), we evaluate prediction performance using two error metrics: absolute percentage error (APE) and mean absolute error (MAE). Given a predicted quantity $Q^{\text{pred}}_{i}$ and the gold quantity $Q^{\text{gold}}_{i}$, the normalized errors are computed as:
$E^{\text{APE}}_{i}=\frac{2}{1+e^{-\alpha\frac{|Q^{\text{pred}}_{i}-Q^{\text{gold}}_{i}|}{Q^{\text{gold}}_{i}+\epsilon}}}-1,$ (8)
$E^{\text{MAE}}_{i}=\frac{2}{1+e^{-\alpha|Q^{\text{pred}}_{i}-Q^{\text{gold}}_{i}|}}-1.$ (9)
Here, $\epsilon$ is a small constant to prevent division by zero, and $\alpha$ is a scaling factor that controls how sharply large errors are penalized, similar to the timeframe prediction evaluation. Confidence calibration is assessed using CRPS, following the same Gaussian assumption as in timeframe prediction. The predicted quantity is modeled as a Gaussian distribution $\mathcal{N}(Q^{\text{pred}}_{i},\sigma^{\text{pred}}_{i})$, and the gold quantity as $\mathcal{N}(Q^{\text{gold}}_{i},\sigma^{\text{gold}}_{i})$. The standard deviations $\sigma^{\text{pred}}_{i}$ and $\sigma^{\text{gold}}_{i}$ are computed using the same formulation as in timeframe prediction.
4 FOReCAst Construction
4.1 Data Source and Question Selection
FOReCAst is constructed from Metaculus,111www.metaculus.com. Examples in Appendix A. an online forecasting platform where forecasters submit probabilistic predictions to questions across various domains. Metaculus aggregates individual forecasts into a continuously updated community prediction, which is finalized just before resolution. Each question has predefined resolution criteria, ensuring verifiable outcomes. To ensure dataset reliability, we include only questions with a definitive resolution and at least 100 forecasts to maintain statistical reliability. Ambiguous or subjectively resolved questions are excluded, and we remove those whose outcomes depend on arbitrary or uncontrollable factors. These steps ensure that FOReCAst consists of high-quality, well-formed forecasting questions with verifiable outcomes.
4.2 Extracting Confidence from Crowdsourced Forecasts
Forecasting distinguishes between physical probabilities—objective likelihoods derived from statistical or scientific models—and human beliefs (Sanders, 1963) about future events. While physical probabilities can be useful, they are often unavailable, particularly for questions involving human behavior, economics, or sociopolitical outcomes. Instead, collective human forecasts offer a more practical confidence estimate, integrating expert reasoning, contextual knowledge, and evolving evidence. For instance, predicting a technological breakthrough depends more on expert assessment and current trends than on rigid probabilistic models. Therefore, confidence in FOReCAst is derived from Metaculus community forecasts, which aggregate predictions from a diverse pool of forecasters. While human predictions are sometimes incorrect, they still serve as a valuable proxy for uncertainty, as they reflect the best available reasoning given the information at the time.
Formal Definition of Gold Confidence.
Gold confidence in FOReCAst is derived from the final Metaculus community prediction before resolution. Instead of directly using the predicted probability for the correct outcome, we compute a log score relative to a uniform baseline, ensuring that confidence reflects how much the forecast deviates from random guessing. This transformation prevents extreme probabilities in inherently uncertain scenarios and makes confidence scores more comparable across different forecasting tasks. The final score is mapped to $(0,1)$ using a sigmoid function.
For Boolean questions, where the human-forecasted probability for the correct outcome is $P^{\text{gold}}$, gold confidence is computed as:
$C^{\text{gold}}=\sigma\left(\frac{\ln P^{\text{gold}}-\ln 0.5}{\ln 2}\right),$ (10)
where $\sigma(x)$ is the sigmoid function.
Similarly, for timeframe prediction and quantity estimation, where the human-forecasted probability density function (PDF) assigns probability to a continuous outcome $x^{\text{gold}}$, gold confidence is computed as:
$C^{\text{gold}}=\sigma\left(\frac{\ln f(x^{\text{gold}})-\ln f_{\text{uniform}}}{2}\right),$ (11)
where $f(x^{\text{gold}})$ is the forecasted probability density at the resolved outcome, and $f_{\text{uniform}}$ is the uniform baseline density over the valid range of values. The denominator 2 ensures numerical stability and scales confidence appropriately.
4.3 Dataset Statistics and Comparison
FOReCAst consists of 2256 forecasting questions, spanning domains such as politics, economics, science, and technology. Each question includes a resolved outcome, a gold confidence score, and a final Metaculus community forecast before resolution. To facilitate model development and evaluation, we split the dataset into 65% training, 10% validation, and 25% test. The full dataset statistics is shown in Appendix B.
Table 2 provides a comparison between FOReCAst and existing forecasting benchmarks. Compared to prior datasets, FOReCAst uniquely emphasizes both forecasting accuracy and confidence calibration, includes a diverse set of forecasting tasks, and is constructed from a well-established crowdsourced platform with rigorous resolution criteria.
Benchmark Question Types Natural Questions Confidence
ForecastQA MCQ
AutoCast Various
ExpTime Boolean
FOReCAst Various
Table 2: Comparison of key features across our benchmark variants, highlighting our evaluation of confidence across different question types.
5 Experiments on FOReCAst
5.1 Experimental Setup
Models.
We evaluate a diverse set of large language models (LLMs) with varying training data cutoffs, model sizes, and instruction tuning. To analyze the impact of knowledge recency, we group models by family and assume the cutoff date is the 1st of the stated month. The models used are shown in Table 3.
Model Family Model Variants Cutoff Date
GPT-2 GPT-2, GPT-2 XL 2017-12-01
Pythia 14M, 160M, 2.8B 2020-03-01
BLOOM 560M, 7B1 2021-12-01
LLaMA LLaMA-7B 2022-08-01
OLMo 1B, 7B, 7B-Instruct 2023-03-01
OLMo-2 7B, 7B-Instruct 2023-12-01
Table 3: Models used in our experiments, grouped by family and ordered by training data cutoff date, including: GPT-2 (Radford et al., 2019), Pythia (Biderman et al., 2023), BLOOM (Scao et al., 2023), LLaMA (Touvron et al., 2023), OLMo (Groeneveld et al., 2024), and OLMo-2 (OLMo et al., 2024).
Inference.
We use 1-shot in-context learning to provide models with a structured example of how to answer forecasting questions. For instruction-tuned models, we add an extra line of instruction to align with their training paradigm. To ensure fair comparison, baseline prompts are kept minimal while maintaining clarity. Confidence is estimated using logit-based normalized heuristics. Full prompts and hyperparameters are provided in Appendix C and  Appendix D.
Model N Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$)
GPT2 401 0.58 0.37 0.42
GPT2-XL 401 0.67 0.32 0.45
Pythia-14m 343 0.59 0.14 0.62
Pythia-160m 343 0.61 0.26 0.53
Pythia-2.8b 343 0.55 0.39 0.44
Bloom-560m 263 0.49 0.41 0.45
Bloom-7b1 263 0.64 0.33 0.32
Llama-7b 226 0.57 0.34 0.52
OLMo-1B 188 0.23 0.16 0.75
OLMo-7B 188 0.21 0.17 0.82
OLMo-7B-Instruct 188 0.65 0.14 0.38
OLMo-2-7B 145 0.51 0.24 0.44
OLMo-2-7B-Instruct 145 0.59 0.30 0.41
Table 4: Performance of forecasting models on Boolean questions in FOReCAst. Reported metrics include the number of evaluated questions (N), accuracy, F1 score (both higher is better), and the Brier score (lower is better).
5.2 Results and Findings
Our experiments on the FOReCAst dataset reveal that forecasting remains highly challenging for current LLMs, particularly in confidence estimation. While models achieve reasonable accuracy in point predictions, their uncertainty estimates—captured by Brier score and CRPS—vary significantly. This suggests that confidence evaluation must be treated as a separate challenge from prediction assessment.
Boolean Questions:
Table 4 shows that while some models achieve reasonable accuracy, their calibration remains inconsistent. For example, GPT2-XL achieves an accuracy of 0.67 but has a relatively high Brier score of 0.45, while OLMo-7B-Instruct achieves similar accuracy (0.65) with a lower Brier score of 0.38, suggesting better confidence estimation. Notably, models with later training cutoffs do not always outperform older ones; for instance, LLaMA-7B has lower accuracy (0.57) and a worse Brier score (0.52) than some earlier models. This indicates that forecasting accuracy depends on multiple factors beyond knowledge recency.
Model N ADE ($\downarrow$) CRPS ($\downarrow$)
GPT2 26 0.99 0.99
GPT2-XL 26 1.00 1.00
Pythia-14m 25 1.00 1.00
Pythia-160m 25 1.00 1.00
Pythia-2.8b 25 0.97 0.96
Bloom-560m 12 1.00 1.00
Bloom-7b1 12 1.00 1.00
Llama-7b 10 0.98 0.98
OLMo-1B 6 1.00 1.00
OLMo-7B 6 1.00 1.00
OLMo-7B-Instruct 6 1.00 1.00
OLMo-2-7B 4 1.00 1.00
OLMo-2-7B-Instruct 4 0.87 0.82
Table 5: Performance of forecasting models on Timeframe Prediction tasks in FOReCAst. Metrics include the number of evaluated questions (N), the normalized absolute days error (ADE), and the Continuous Ranked Probability Score (CRPS), where lower values indicate better performance.
Model N APE ($\downarrow$) MAE ($\downarrow$) CRPS ($\downarrow$)
GPT2 81 0.23 0.87 0.86
GPT2-XL 81 0.03 0.75 0.72
Pythia-14m 77 0.17 0.88 0.85
Pythia-160m 77 0.03 0.78 0.76
Pythia-2.8b 77 0.07 0.81 0.79
Bloom-560m 43 0.05 0.73 0.71
Bloom-7b1 43 0.06 0.75 0.72
Llama-7b 32 0.06 0.64 0.61
OLMo-1B 23 0.21 0.79 0.76
OLMo-7B 23 0.22 0.84 0.82
OLMo-7B-Instruct 23 0.24 0.80 0.77
OLMo-2-7B 20 0.02 0.65 0.62
OLMo-2-7B-Instruct 20 0.09 0.60 0.57
Table 6: Performance of forecasting models on Quantity Estimation tasks in FOReCAst. Metrics include the number of evaluated questions (N), normalized absolute percentage error (APE), mean absolute error (MAE), and Continuous Ranked Probability Score (CRPS), where lower values are better.
Timeframe Prediction:
Table 5 reveals that most models struggle with predicting event timing, as both ADE and CRPS remain near the worst-case scenario of 1.0. Across all models, only one achieves a CRPS below 0.9: OLMo-2-7B-Instruct (CRPS = 0.82), which still indicates substantial uncertainty. Even models with more recent training data, such as OLMo-2-7B (CRPS = 0.99), fail to make well-calibrated temporal forecasts. These results suggest that even when models correctly anticipate whether an event will happen, quantifying when it will occur remains a major challenge.
Quantity Estimation:
Table 6 highlights that models differ widely in their ability to estimate numerical values. Some achieve relatively low APE, such as GPT2-XL (APE = 0.03) and Pythia-160M (APE = 0.02), yet their confidence calibration, CRPS, does not always align with their point prediction performance. For instance, GPT2-XL has a CRPS of 0.72, while Pythia-160M has a slightly higher CRPS of 0.76, despite achieving lower APE. In contrast, OLMo-2-7B-Instruct, which has a relatively higher APE of 0.08, achieves the lowest CRPS (0.57) among all models. These results indicate that prediction quality and confidence calibration do not necessarily improve together, reinforcing the complexity of numerical forecasting.
Impact of Model Size on Forecasting Performance
Larger models do not consistently improve forecasting performance. Within the Pythia family, Pythia-2.8b shows occasional gains in accuracy and calibration over Pythia-14m and Pythia-160m, but the improvements are not uniform across all metrics. Similarly, while OLMo-2-7B variants sometimes achieve lower CRPS and MAE in quantity estimation, these benefits often come with trade-offs in point prediction. These results suggest that model size alone is not a reliable predictor of forecasting performance. This highlights the importance of developing task-specific techniques rather than relying on ever-larger models to solve the forecasting problem.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
OLMo-7B 0.21 0.17 0.82 1.00 1.00 0.22 0.84 0.82
OLMo-7B-Instruct 0.65 0.14 0.38 1.00 1.00 0.24 0.80 0.77
OLMo-2-7B 0.51 0.24 0.44 1.00 1.00 0.02 0.65 0.62
OLMo-2-7B-Instruct 0.59 0.30 0.41 0.87 0.82 0.09 0.60 0.57
Table 7: Effect of instruction tuning on forecasting performance across boolean, timeframe, and quantity estimation questions. Metrics include accuracy, F1, and Brier score for binary questions; ADE and CRPS (T) for timeframe prediction; and APE, MAE, and CRPS (Q) for quantity estimation.
Impact of Instruction Tuning on Forecasting Performance
Table 7 compares base and instruct-tuned variants of OLMo-7B and OLMo-2-7B across Boolean, timeframe, and quantity forecasting tasks. For Boolean questions, OLMo-7B-Instruct achieves higher accuracy (0.65 vs. 0.21) and a lower Brier score (0.38 vs. 0.82), indicating better confidence calibration. In timeframe prediction, the instruct-tuned OLMo-2-7B-Instruct improves uncertainty estimation, with an ADE of 0.87 and CRPS of 0.82, compared to 0.9981 and 0.9970 for the base OLMo-2-7B. For quantity estimation, instruct-tuned models have slightly higher APE but lower MAE and CRPS, suggesting better uncertainty calibration. These results indicate that instruction tuning enhances confidence estimation, even if it does not always improve point prediction accuracy. This suggests a trade-off where instruct-tuned models prioritize more reliable uncertainty quantification.
Aggregation Method N Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$)
Majority Vote 226 0.56 0.24 0.53
Highest Confidence 226 0.57 0.34 0.52
Weighted Average 226 0.56 0.24 0.29
Logit Mean Probability 226 0.56 0.24 0.62
Bayesian Aggregation 226 0.58 0.35 0.55
Table 8: Ablation study on different aggregation methods for extracting predictions and confidence from Llama-7B outputs.
Impact of Aggregation Methods on Forecasting Performance
Table 8 compares different aggregation methods for deriving the final prediction and confidence estimate from the top 10 outputs of Llama-7B. Bayesian Aggregation achieves the highest accuracy (0.5796) and F1 score (0.3485), suggesting it is the most effective at identifying correct outcomes. However, Weighted Average yields a significantly lower Brier score (0.2914), indicating superior confidence calibration compared to other methods.
Majority Vote, Highest Confidence, and Logit Mean Probability produce comparable accuracy and F1 scores but have noticeably higher Brier scores, suggesting weaker uncertainty estimation. These results highlight that even when point prediction performance is similar, aggregation methods substantially impact confidence reliability. The challenge remains in developing techniques that optimize both accuracy and calibration simultaneously, emphasizing the importance of uncertainty-aware forecasting.
6 Related Work
Recent forecasting benchmarks focus on event prediction but largely overlook confidence calibration. OpenForecast (Wang et al., 2025) introduces a large-scale dataset for open-ended, multi-step event forecasting but does not assess model confidence. ForecastBench (Karger et al., 2024) evaluates binary (Yes/No) forecasting by prompting models for direct probability estimates, but since it queries each option separately, the assigned probabilities do not necessarily sum to 1, leading to potential inconsistencies. Neither benchmark systematically evaluates confidence calibration, a crucial aspect for reliable forecasting in real-world applications.
Beyond forecasting, several benchmarks assess language models’ reasoning and inference capabilities. COPA (Roemmele et al., 2011) evaluates causal reasoning by presenting a premise and two alternatives, requiring models to select the more plausible cause or effect. HellaSwag Zellers et al. (2019) challenges models with sentence completion tasks that demand commonsense reasoning, where models must choose the most sensible continuation of a given scenario. PRobELM (Yuan et al., 2024b) assesses models’ capacity to rank scenarios by plausibility, bridging the gap between factual accuracy and world knowledge. While these benchmarks provide insights into models’ reasoning abilities, they do not address the challenges of forecasting future events.
7 Conclusion
We introduce FOReCAst, a benchmark for evaluating both forecasting accuracy and confidence calibration in language models. Unlike existing datasets, FOReCAst explicitly assesses confidence alongside predictions. Our results show that current models struggle with both prediction and well-calibrated confidence, underscoring the need for improved uncertainty estimation and confidence calibration.
Limitations
While FOReCAst represents a significant step toward evaluating forecasting accuracy and confidence calibration, there are inherent limitations that we view as opportunities for future research rather than fundamental shortcomings of our work. First, our dataset is constructed solely from Metaculus, which may not fully represent the global diversity of forecasting practices or question domains. Second, our method for deriving gold confidence relies on community forecasts and heuristic transformations that might not capture all nuances of human uncertainty. Lastly, our focus on English-language forecasts limits the benchmark’s applicability across different languages and cultural contexts. Addressing these issues is part of our future work agenda.
Ethical Statement
FOReCAst is built from data sourced exclusively from Metaculus, an English-language forecasting platform. As a result, the dataset may embody the linguistic, cultural, and social biases inherent in its user community. These biases could affect both question selection and confidence judgments. We acknowledge these concerns and stress that our benchmark is intended as an initial step toward more inclusive forecasting evaluations. Future efforts should aim to incorporate data from a broader range of platforms and languages to mitigate these biases.
Acknowledgements
All authors of this paper are supported by the ERC grant AVeriTeC (GA 865958). Andreas Vlachos receives further support from the DARPA program SciFy.
Appendix A Example Metaculus Questions
To illustrate how human forecasts evolve over time, we present two questions from different domains: Q1 is in the business and geopolitics domain and Q2 is in the technology domain.
Q1: Will TikTok become available in the US on both the App Store and Google Play before April 5, 2025?
Q2: When will a SpaceX Starship reach orbit?
For Q1, Figure 1 shows how community forecasts changed over time, while Figure 2 presents the histogram of the final forecast distribution.
Figure 1: Community prediction trend for a Metaculus question on TikTok’s availability in the US.
Figure 2: Histogram of final community forecasts.
For Q2, Figure 3 tracks forecast updates, while Figure 4 shows the final probability density function (PDF) of predicted launch dates.
Figure 3: Community prediction trend for SpaceX Starship’s first orbital launch.
Figure 4: Probability density function of final community forecasts for SpaceX Starship reaching orbit.
Appendix B Dataset Statistics
Split Boolean Timeframe Quantity Total
Training 1142 90 223 1465
Validation 175 13 35 223
Test 441 36 91 568
Total 1758 139 349 2256
Table 9: Dataset statistics for FOReCAst, showing the distribution of questions across different forecasting types, with the overall total in the last column.
Table 9 presents detailed dataset statistics, including the total number of questions and their distribution across Boolean Questions, Timeframe Prediction, and Quantity Estimation tasks.
Appendix C Prompts
To ensure a fair and consistent evaluation across models, we use simple one-shot prompts with structured outputs in JSON format. For instruction-tuned models, we provide an additional instruction line specifying the task. The prompts are designed for three forecasting question types: Boolean Questions (Yes/No), Timeframe Prediction (YYYY-MM-DD), and Quantity Estimation (numeric values).
C.1 Instruction-Tuned Models
For models containing "Instruct" in their name, we use the following prompts:
Quantity Estimation
You are an AI assistant providing precise numerical forecasts.
Answer the following question with a single numeric value in JSON format.
Example:
Q: How much global photovoltaic energy generation was deployed
   by the end of 2020?
A: { "value": 738 }
Q: \$question
A: { "value": "
Timeframe Prediction
You are an AI assistant providing precise date forecasts.
Answer the following question with a single date in YYYY-MM-DD format in JSON.
Example:
Q: When did an AI system achieve a significant victory against
   a professional human in Starcraft 2?
A: { "value": "2019-01-24" }
Q: \$question
A: { "value": "
Boolean Questions
You are an AI assistant providing binary (Yes/No) answers.
Answer the following question with "Yes" or "No" in JSON format.
Example:
Q: Will we confirm evidence for megastructures orbiting the
   star KIC 8462852?
A: { "value": "No" }
Q: \$question
A: { "value": "
C.2 Base Models
For non-instruction-tuned models, we use the same examples but without additional instructions:
Quantity Estimation
Q: How much global photovoltaic energy generation was deployed
   by the end of 2020?
A: { "value": 738 }
Q: \$question
A: { "value": "
Timeframe Prediction
Q: When did an AI system achieve a significant victory against
   a professional human in Starcraft 2?
A: { "value": "2019-01-24" }
Q: \$question
A: { "value": "
Boolean Questions
Q: Will we confirm evidence for megastructures orbiting the
   star KIC 8462852?
A: { "value": "No" }
Q: \$question
A: { "value": "
Appendix D Hyperparameter Settings
D.1 Generation Hyperparameters
We generate responses using temperature-based sampling with the following hyperparameters:
  • max_length = 200
  • do_sample = True
  • top_k = 50
  • top_p = 0.9
Among the generated outputs, we select the one with the highest confidence as the final prediction. All experiments are conducted using full precision on an NVIDIA RTX 8000 GPU.
D.2 Evaluation Hyperparameters
The scaling factor $\alpha$ in Equation 4, Equation 8, and Equation 9 is set to 0.05. For Equation 5 and Equation 6, we set $\sigma_{\max}$ to 30 and $\sigma_{\min}$ to 1 for Timeframe Prediction, while for Quantity Estimation, $\sigma_{\max}$ is 20 and $\sigma_{\min}$ is 1. These values ensure that evaluation metrics appropriately scale errors and confidence calibration.
Appendix E Additional Results
This section provides extended results categorized by the training data cutoff date of each model. Forecasting performance depends on model architecture, scale, and knowledge recency, so we evaluate models with different cutoff dates to examine how access to more recent information influences prediction accuracy and confidence calibration.
Models trained after certain event resolutions may have indirectly encountered outcome-related information, potentially affecting evaluation fairness. This should be considered when interpreting results.
Detailed model-specific performance metrics for Boolean Questions, Timeframe Prediction, and Quantity Estimation are presented in Table 10 to Table 15.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
GPT2 0.58 0.37 0.42 0.99 0.99 0.23 0.87 0.86
GPT2-XL 0.67 0.32 0.45 1.00 1.00 0.03 0.75 0.72
Pythia-14m 0.60 0.15 0.62 1.00 1.00 0.18 0.86 0.84
Pythia-160m 0.61 0.25 0.53 1.00 1.00 0.03 0.76 0.74
Pythia-2.8b 0.53 0.37 0.46 0.97 0.96 0.08 0.79 0.77
Bloom-560m 0.48 0.42 0.48 1.00 1.00 0.06 0.77 0.75
Bloom-7b1 0.63 0.32 0.35 0.97 0.95 0.05 0.76 0.74
Llama-7b 0.55 0.37 0.53 0.94 0.93 0.08 0.72 0.70
OLMo-1B 0.23 0.20 0.79 1.00 1.00 0.15 0.84 0.82
OLMo-7B 0.21 0.17 0.80 1.00 1.00 0.27 0.88 0.87
OLMo-7B-Instruct 0.67 0.33 0.39 0.92 0.91 0.29 0.80 0.78
OLMo-2-7B 0.54 0.33 0.50 0.92 0.91 0.04 0.70 0.68
OLMo-2-7B-Instruct 0.59 0.42 0.40 0.84 0.83 0.13 0.69 0.67
Table 10: Combined forecasting performance for cutoff 2017-12-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
GPT2 0.59 0.38 0.41 0.99 0.99 0.23 0.89 0.87
GPT2-XL 0.67 0.32 0.45 1.00 1.00 0.03 0.77 0.74
Pythia-14m 0.59 0.14 0.62 1.00 1.00 0.17 0.88 0.85
Pythia-160m 0.61 0.26 0.53 1.00 1.00 0.03 0.78 0.76
Pythia-2.8b 0.55 0.39 0.44 0.97 0.96 0.07 0.81 0.79
Bloom-560m 0.50 0.44 0.46 1.00 1.00 0.05 0.78 0.77
Bloom-7b1 0.63 0.33 0.35 0.97 0.95 0.05 0.79 0.76
Llama-7b 0.55 0.37 0.53 0.94 0.93 0.06 0.72 0.70
OLMo-1B 0.23 0.19 0.78 1.00 1.00 0.16 0.87 0.85
OLMo-7B 0.20 0.15 0.80 1.00 1.00 0.26 0.89 0.87
OLMo-7B-Instruct 0.66 0.30 0.40 0.92 0.91 0.27 0.80 0.78
OLMo-2-7B 0.53 0.33 0.51 0.92 0.90 0.03 0.71 0.68
OLMo-2-7B-Instruct 0.59 0.41 0.40 0.83 0.82 0.12 0.70 0.68
Table 11: Combined forecasting performance for cutoff 2020-03-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
GPT2 0.61 0.35 0.38 1.00 1.00 0.26 0.85 0.84
GPT2-XL 0.67 0.26 0.45 1.00 1.00 0.03 0.74 0.70
Pythia-14m 0.60 0.13 0.65 1.00 1.00 0.22 0.86 0.84
Pythia-160m 0.63 0.24 0.52 1.00 1.00 0.02 0.73 0.70
Pythia-2.8b 0.52 0.33 0.44 1.00 1.00 0.07 0.77 0.74
Bloom-560m 0.50 0.44 0.45 1.00 1.00 0.05 0.73 0.71
Bloom-7b1 0.63 0.33 0.32 1.00 1.00 0.06 0.75 0.72
Llama-7b 0.57 0.36 0.52 0.99 0.98 0.05 0.65 0.62
OLMo-1B 0.24 0.18 0.77 1.00 1.00 0.22 0.84 0.82
OLMo-7B 0.22 0.18 0.80 1.00 1.00 0.23 0.85 0.84
OLMo-7B-Instruct 0.67 0.18 0.40 0.92 0.92 0.22 0.76 0.74
OLMo-2-7B 0.53 0.27 0.49 0.93 0.92 0.03 0.67 0.64
OLMo-2-7B-Instruct 0.61 0.38 0.39 0.86 0.84 0.10 0.64 0.61
Table 12: Combined forecasting performance for cutoff 2021-12-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
GPT2 0.62 0.36 0.36 1.00 1.00 0.26 0.81 0.80
GPT2-XL 0.68 0.27 0.44 1.00 1.00 0.03 0.74 0.70
Pythia-14m 0.61 0.16 0.65 1.00 1.00 0.19 0.87 0.85
Pythia-160m 0.65 0.25 0.53 1.00 1.00 0.02 0.73 0.70
Pythia-2.8b 0.52 0.32 0.45 1.00 1.00 0.08 0.78 0.74
Bloom-560m 0.50 0.40 0.43 1.00 1.00 0.05 0.72 0.71
Bloom-7b1 0.65 0.33 0.31 1.00 1.00 0.08 0.78 0.74
Llama-7b 0.57 0.34 0.52 0.98 0.98 0.06 0.64 0.61
OLMo-1B 0.23 0.18 0.78 1.00 1.00 0.22 0.83 0.80
OLMo-7B 0.23 0.17 0.82 1.00 1.00 0.22 0.87 0.86
OLMo-7B-Instruct 0.65 0.12 0.40 0.91 0.90 0.21 0.78 0.75
OLMo-2-7B 0.54 0.26 0.47 0.91 0.91 0.03 0.68 0.65
OLMo-2-7B-Instruct 0.60 0.33 0.40 0.93 0.90 0.10 0.65 0.62
Table 13: Combined forecasting performance for cutoff 2022-08-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
GPT2 0.60 0.36 0.38 1.00 1.00 0.26 0.74 0.72
GPT2-XL 0.65 0.24 0.42 1.00 1.00 0.04 0.71 0.65
Pythia-14m 0.60 0.16 0.64 1.00 1.00 0.22 0.85 0.82
Pythia-160m 0.63 0.26 0.50 1.00 1.00 0.02 0.69 0.66
Pythia-2.8b 0.52 0.34 0.44 1.00 1.00 0.11 0.76 0.72
Bloom-560m 0.53 0.42 0.41 1.00 1.00 0.07 0.70 0.67
Bloom-7b1 0.66 0.39 0.30 1.00 1.00 0.10 0.79 0.76
Llama-7b 0.55 0.32 0.51 1.00 1.00 0.06 0.60 0.57
OLMo-1B 0.23 0.16 0.75 1.00 1.00 0.21 0.79 0.76
OLMo-7B 0.21 0.17 0.82 1.00 1.00 0.22 0.84 0.82
OLMo-7B-Instruct 0.65 0.14 0.38 1.00 1.00 0.24 0.80 0.77
OLMo-2-7B 0.51 0.24 0.44 1.00 1.00 0.02 0.65 0.62
OLMo-2-7B-Instruct 0.59 0.33 0.41 0.91 0.87 0.12 0.61 0.57
Table 14: Combined forecasting performance for cutoff 2023-03-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
GPT2 0.59 0.31 0.37 1.00 1.00 0.25 0.73 0.72
GPT2-XL 0.71 0.30 0.42 1.00 1.00 0.04 0.69 0.64
Pythia-14m 0.62 0.14 0.65 1.00 1.00 0.25 0.85 0.83
Pythia-160m 0.67 0.31 0.51 1.00 1.00 0.02 0.67 0.65
Pythia-2.8b 0.52 0.29 0.45 1.00 1.00 0.12 0.73 0.69
Bloom-560m 0.52 0.38 0.41 1.00 1.00 0.03 0.67 0.65
Bloom-7b1 0.68 0.36 0.30 1.00 1.00 0.11 0.78 0.75
Llama-7b 0.54 0.27 0.55 0.96 0.94 0.06 0.62 0.59
OLMo-1B 0.23 0.21 0.75 1.00 1.00 0.22 0.76 0.74
OLMo-7B 0.22 0.23 0.83 1.00 1.00 0.22 0.82 0.81
OLMo-7B-Instruct 0.68 0.16 0.36 1.00 1.00 0.27 0.81 0.78
OLMo-2-7B 0.51 0.24 0.44 1.00 1.00 0.02 0.65 0.62
OLMo-2-7B-Instruct 0.59 0.30 0.41 0.87 0.82 0.09 0.60 0.57
Table 15: Combined forecasting performance for cutoff 2023-12-01. CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
These results highlight trends in forecasting accuracy and confidence calibration across models with different knowledge recency.

### v3

@@
-The Future Outcome Reasoning and Confidence Assessment Benchmark
+Introducing FOReCAst: The Future Outcome Reasoning and Confidence Assessment Benchmark
 Abstract
@@
 However existing forecasting benchmarks largely lack comprehensive confidence assessment, focus on limited question types, and often consist of artificial questions that do not align with real-world human forecasting needs.
-To address these gaps, we introduce FOReCAst (Future Outcome Reasoning and Confidence Assessment), a benchmark that evaluates models’ ability to make predictions and their confidence in them.
+To address these gaps, we introduce FOReCAst (Future Outcome Reasoning and Confidence Assessment), [^1] Publicly availble at https://huggingface.co/datasets/MoyYuan/FOReCAst a benchmark that evaluates models’ ability to make predictions and their confidence in them.
 FOReCAst spans diverse forecasting scenarios involving Boolean questions, timeframe prediction, and quantity estimation, enabling a comprehensive evaluation of both prediction accuracy and confidence calibration for real-world applications.
-The Future Outcome Reasoning and Confidence Assessment Benchmark
-Zhangdie Yuan University of Cambridge zy317@cam.ac.uk                        Zifeng Ding University of Cambridge zd320@cam.ac.uk                        Andreas Vlachos University of Cambridge av308@cam.ac.uk
+1 Introduction
+Recent advances in large language models (LLMs) have significantly improved their performance across a wide range of natural language processing (NLP) tasks.
+Alongside these developments, various benchmarks and datasets have been introduced to effectively assess the capabilities of LLMs, particularly in terms of knowledge and reasoning (Zellers et al., 2019; Guo et al., 2023).
+Fact-based benchmarks, such as TruthfulQA (Lin et al., 2022), evaluate LLMs based on factual correctness, focusing on tasks like retrieving and verifying facts that are already known.
+Forecasting is a crucial yet challenging task across various domains, including technology, economics, and public policy.
+Unlike tasks that rely on retrieving and verifying existing knowledge, forecasting requires predicting plausible outcomes for future events, often under uncertainty and incomplete information.
+This makes forecasting particularly difficult, as models must infer trends, assess probabilities, and adapt to new information.
+Several datasets have been introduced to evaluate LLMs’ forecasting capabilities.
+ForecastQA (Jin et al., 2020) uses a multiple-choice format where models predict future outcomes, but it lacks confidence assessment.
+AutoCast (Zou et al., 2022) incorporates confidence intervals; however, its confidence estimates are not specifically designed for forecasting tasks.
+Other datasets, such as ExpTime (Yuan et al., 2024a), are artificially generated from structured data and focus on explainable event forecasting based on temporal knowledge graphs.
+All of these aforementioned benchmarks overlook a crucial aspect of forecasting: confidence evaluation.
+Confidence plays a central role in forecasting, as predictions about unresolved events inherently lack definitive correctness at the time of evaluation.
+Predictions made with absolute certainty are undesirable, even if they ultimately prove to be correct, because they fail to account for the uncertain nature of future events.
+Moreover, miscalibrated confidence can lead to poor decision-making: overconfident yet incorrect forecasts may result in costly errors, while underconfident but accurate predictions can erode trust in the model.
+Therefore, well-calibrated confidence scores are as crucial as the accuracy of the predictions themselves.
+To address these gaps, we present FOReCAst: Future Outcome Reasoning and Confidence Assessment.
+FOReCAst focuses on three distinct types of forecasting questions, shown in Table 1: (1) Boolean Questions, such as "Will there be a Frontier AI lab in China before 2026?"; (2) Timeframe Prediction, such as "When will OpenAI announce GPT-5?"; and (3) Quantity Estimation, such as "How many spacecrafts will land on the moon in each of the following years?" We conduct experiments using a range of models differing in size, training objectives, and training cutoff times, and explore multiple methods for estimating model confidence.
+Our results reveal that forecasting remains highly challenging for current LLMs, particularly in confidence evaluation, with no consistent correlation between prediction accuracy and calibration.
+While larger models sometimes improve performance, the effect is inconsistent and highly task-dependent.
 Type Question Resolution Confidence
@@
 Table 1: Examples of forecasting questions with their resolutions and confidence scores.
-1 Introduction
-Recent advances in large language models (LLMs) have significantly improved their performance across a wide range of natural language processing (NLP) tasks.
-Alongside these developments, various benchmarks and datasets have been introduced to effectively assess the capabilities of LLMs, particularly in terms of knowledge and reasoning (Zellers et al., 2019; Guo et al., 2023).
-Fact-based benchmarks, like TruthfulQA (Lin et al., 2022) evaluate LLMs based on factual correctness, focusing on tasks like retrieving and verifying facts that are known.
-Forecasting is a crucial yet challenging task across various domains, including technology, economics, and public policy.
-Unlike tasks that rely on retrieving and verifying existing knowledge, forecasting requires predicting plausible outcomes for future events, often under uncertainty and incomplete information.
-This makes forecasting particularly difficult, as models must infer trends, assess probabilities, and adapt to new information.
-Several datasets have been introduced to evaluate LLMs’ forecasting capabilities.
-ForecastQA (Jin et al., 2020) uses a multiple-choice format where models predict future outcomes, but it lacks confidence assessment.
-AutoCast (Zou et al., 2022) incorporates confidence intervals, however its confidence estimates are not designed for forecasting.
-Other datasets such as ExpTime (Yuan et al., 2024a) are artificially generated from structured data, focusing on explainable event forecasting based on temporal knowledge graphs.
-All of these aforementioned benchmarks ignore a crucial aspect of forecasting: confidence evaluation.
-Confidence plays a central role in forecasting, as predictions about unresolved events inherently lack definitive correctness at the time of evaluation.
-Predictions made with absolute certainty are undesirable, even if they ultimately prove to be correct, because they fail to account for the uncertain nature of future events.
-Moreover, miscalibrated confidence can lead to poor decision-making: overconfident yet incorrect forecasts may result in costly errors, while underconfident but accurate predictions can erode trust in the model.
-Therefore, well-calibrated confidence scores are as crucial as the accuracy of the predictions themselves.
-To address these gaps, we present FOReCAst: Future Outcome Reasoning and Confidence Assessment.
-FOReCAst focuses on three distinct types of forecasting questions, shown in Table 1: (1) Boolean questions, such as "Will there be a Frontier AI lab in China before 2026?"; (2) Timeframe Prediction, such as "When will OpenAI announce GPT-5?"; and (3) Quantity Estimation, such as "How many spacecrafts will land on the moon in each of the following years?" We conduct experiments using a range of models differing in size, training objectives, and cutoff times, and explore multiple methods for estimating model confidence.
-Our results reveal that forecasting remains highly challenging for current LLMs, particularly in confidence evaluation, with no direct correlation between prediction performance and confidence calibration, and while larger models sometimes improve performance, the effect is inconsistent.
 2 FOReCAst: Problem Formulation
-System responses in FOReCAst consist of (1) a prediction answering a question given the available information and (2) a confidence score in the prediction. This
-ensures a comprehensive assessment of forecasting, accounting for both correctness and confidence calibration.
+System responses in FOReCAst consist of (1) a prediction answering a question given the available information and (2) a confidence score in the prediction.
+This ensures a comprehensive assessment of forecasting, accounting for both correctness and confidence calibration.
 Questions belong to three types.
@@
 The objective is for $M$ to produce an answer $A$ in $\mathcal{A}$ and an associated confidence score $C$, where:
-$M(Q)\to(A,C),\quad C\in[0,1].$ (1)
-$A=\arg\max_{a\in\mathcal{A}}P(X=a|Q,\mathcal{K}(t_{train})).$ (2)
+$M(Q)\to(A,C)\quad where\quad A=\arg\max_{a\in\mathcal{A}}P(X=a|Q,\mathcal{K}(t_{train}))\quad and\quad C\in[0,1].$ (1)
 Here, $\mathcal{K}(t_{train})$ represents the knowledge accessible to the model up to time $t_{train}$.
@@
 For Boolean questions, where the answer space is $\mathcal{A}=\{\text{Yes},\text{No}\}$, prediction performance is evaluated using standard classification metrics, including accuracy and F1-score.
-Confidence calibration is assessed using a modified version of the Brier score (Brier, 1950), which measures the mean squared error between predicted confidence and the gold confidence provided in the dataset, which we assume is provided, and
-represents the likelihood of an event occurring. The modified Brier score is defined as:
+Confidence calibration is assessed using a modified version of the Brier score (Brier, 1950), which measures the mean squared error between predicted confidence and the gold confidence provided in the dataset, which we assume is provided, and represents the likelihood of an event occurring.
+The modified Brier score is defined as:
 $\text{Brier}=\frac{1}{N}\sum_{i=1}^{N}(C^{\text{pred}}_{i}-C^{\text{gold}}_{i})^{2},$ (2)
@@
 Specifically, we assume that both the predicted and the gold confidence predictions follow Gaussian distributions, namely $\mathcal{N}(D^{\text{pred}}_{i},\sigma^{\text{pred}}_{i})$ and $\mathcal{N}(D^{\text{gold}}_{i},\sigma^{\text{gold}}_{i})$ respectively, where the standard deviations are computed as:
-$\sigma^{\text{pred}}_{i}=\sigma_{\max}\cdot(1-C^{\text{pred}}_{i})+\sigma_{\min}\cdot C^{\text{pred}}_{i},$ (5)
-$\sigma^{\text{gold}}_{i}=\sigma_{\max}\cdot(1-C^{\text{gold}}_{i})+\sigma_{\min}\cdot C^{\text{gold}}_{i}.$ (6)
+$\sigma^{\text{pred}}_{i}=\sigma_{\max}\cdot(1-C^{\text{pred}}_{i})+\sigma_{\min}\cdot C^{\text{pred}}_{i},\quad\sigma^{\text{gold}}_{i}=\sigma_{\max}\cdot(1-C^{\text{gold}}_{i})+\sigma_{\min}\cdot C^{\text{gold}}_{i}.$ (4)
 Here, $C^{\text{pred}}_{i}$ is the model’s predicted confidence for the $i$th question, and $C^{\text{gold}}_{i}$ is the corresponding gold confidence provided in our dataset.
@@
 Given a predicted quantity $Q^{\text{pred}}_{i}$ and the gold quantity $Q^{\text{gold}}_{i}$, the normalized errors are computed as:
-$E^{\text{APE}}_{i}=\frac{2}{1+e^{-\alpha\frac{|Q^{\text{pred}}_{i}-Q^{\text{gold}}_{i}|}{Q^{\text{gold}}_{i}+\epsilon}}}-1,$ (8)
-$E^{\text{MAE}}_{i}=\frac{2}{1+e^{-\alpha|Q^{\text{pred}}_{i}-Q^{\text{gold}}_{i}|}}-1.$ (9)
+$E^{\text{APE}}_{i}=\frac{2}{1+e^{-\alpha\frac{|Q^{\text{pred}}_{i}-Q^{\text{gold}}_{i}|}{Q^{\text{gold}}_{i}+\epsilon}}}-1,\quad E^{\text{MAE}}_{i}=\frac{2}{1+e^{-\alpha|Q^{\text{pred}}_{i}-Q^{\text{gold}}_{i}|}}-1.$ (6)
 Here, $\epsilon$ is a small constant to prevent division by zero, and $\alpha$ is a scaling factor that controls how sharply large errors are penalized, similar to the timeframe prediction evaluation.
@@
 4.1 Data Source and Question Selection
-FOReCAst is constructed from Metaculus,111www.metaculus.com.
+FOReCAst is constructed from Metaculus,222www.metaculus.com.
 Examples in Appendix A.
@@
 These steps ensure that FOReCAst consists of high-quality, well-formed forecasting questions with verifiable outcomes.
+The dataset spans a diverse range of domains such as politics, economics, health, and technology, with questions distributed across a wide temporal range; detailed statistics are provided in Table 9.
 4.2 Extracting Confidence from Crowdsourced Forecasts
@@
 While human predictions are sometimes incorrect, they still serve as a valuable proxy for uncertainty, as they reflect the best available reasoning given the information at the time.
-Formal Definition of Gold Confidence.
-Gold confidence in FOReCAst is derived from the final Metaculus community prediction before resolution.
-Instead of directly using the predicted probability for the correct outcome, we compute a log score relative to a uniform baseline, ensuring that confidence reflects how much the forecast deviates from random guessing.
-This transformation prevents extreme probabilities in inherently uncertain scenarios and makes confidence scores more comparable across different forecasting tasks.
-The final score is mapped to $(0,1)$ using a sigmoid function.
+Formal Definition of Gold Confidence
+Gold confidence in FOReCAst is derived from the final Metaculus community prediction just before question resolution.
+Rather than directly using the predicted probability for the correct outcome, we compute a log score relative to a uniform baseline, following Metaculus’s established approach for evaluating probabilistic forecasts  [^3] https://www.metaculus.com/help/scores-faq/.
+This transformation ensures that confidence reflects not only correctness, but also the amount of informational gain over random guessing.
+It provides a consistent interpretation of confidence across tasks and scales, particularly when outcomes lie in continuous spaces.
 For Boolean questions, where the human-forecasted probability for the correct outcome is $P^{\text{gold}}$, gold confidence is computed as:
 $C^{\text{gold}}=\sigma\left(\frac{\ln P^{\text{gold}}-\ln 0.5}{\ln 2}\right),$ (7)
-where $\sigma(x)$ is the sigmoid function.
-Similarly, for timeframe prediction and quantity estimation, where the human-forecasted probability density function (PDF) assigns probability to a continuous outcome $x^{\text{gold}}$, gold confidence is computed as:
+where $\sigma(x)$ is the sigmoid function. This yields a score of 0.5 for a random-guess forecast ($P^{\text{gold}}=0.5$), and smoothly increases or decreases with higher or lower confidence.
+For timeframe prediction and quantity estimation, where the human-forecasted probability density function (PDF) assigns likelihood to a continuous outcome $x^{\text{gold}}$, gold confidence is computed as:
 $C^{\text{gold}}=\sigma\left(\frac{\ln f(x^{\text{gold}})-\ln f_{\text{uniform}}}{2}\right),$ (8)
-where $f(x^{\text{gold}})$ is the forecasted probability density at the resolved outcome, and $f_{\text{uniform}}$ is the uniform baseline density over the valid range of values.
-The denominator 2 ensures numerical stability and scales confidence appropriately.
+where $f(x^{\text{gold}})$ is the forecasted probability density at the resolved outcome, and $f_{\text{uniform}}$ is the uniform baseline density over the valid outcome range.
+The denominator of 2 stabilizes the scale and avoids over-amplifying minor density differences.
+This approach accounts for the resolution granularity of each question and ensures confidence is not inflated in uncertain or diffuse distributions.
+While this scoring formulation is tailored to Metaculus, it can be adapted or replaced if used with other forecasting platforms or scoring protocols.
+Our goal is to align confidence with meaningful deviations from uncertainty, while maintaining a consistent and interpretable signal across all tasks.
 4.3 Dataset Statistics and Comparison
@@
 Benchmark Question Types Natural Questions Confidence
-ForecastQA MCQ
-AutoCast Various
-ExpTime Boolean
+ForecastQA (Jin et al., 2020) MCQ
+AutoCast (Zou et al., 2022) Various
+ExpTime (Yuan et al., 2024a) Boolean
 FOReCAst Various
@@
 Model Family Model Variants Cutoff Date
-GPT-2 GPT-2, GPT-2 XL 2017-12-01
-Pythia 14M, 160M, 2.8B 2020-03-01
-BLOOM 560M, 7B1 2021-12-01
-LLaMA LLaMA-7B 2022-08-01
-OLMo 1B, 7B, 7B-Instruct 2023-03-01
-OLMo-2 7B, 7B-Instruct 2023-12-01
-Table 3: Models used in our experiments, grouped by family and ordered by training data cutoff date, including: GPT-2 (Radford et al., 2019), Pythia (Biderman et al., 2023), BLOOM (Scao et al., 2023), LLaMA (Touvron et al., 2023), OLMo (Groeneveld et al., 2024), and OLMo-2 (OLMo et al., 2024).
+GPT-2 (Radford et al., 2019) GPT-2, GPT-2 XL 2017-12-01
+Pythia (Biderman et al., 2023) 14M, 160M, 2.8B 2020-03-01
+BLOOM (Scao et al., 2023) 560M, 7B1 2021-12-01
+LLaMA (Touvron et al., 2023) LLaMA-7B 2022-08-01
+OLMo (Groeneveld et al., 2024) 1B, 7B, 7B-Instruct 2023-03-01
+OLMo-2 (OLMo et al., 2024) 7B, 7B-Instruct 2023-12-01
+Table 3: Models used in our experiments, grouped by family and ordered by training data cutoff date.
 Inference.
@@
 Full prompts and hyperparameters are provided in Appendix C and  Appendix D.
+[…]
 Model N Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$)
@@
 Reported metrics include the number of evaluated questions (N), accuracy, F1 score (both higher is better), and the Brier score (lower is better).
-5.2 Results and Findings
-Our experiments on the FOReCAst dataset reveal that forecasting remains highly challenging for current LLMs, particularly in confidence estimation.
-While models achieve reasonable accuracy in point predictions, their uncertainty estimates—captured by Brier score and CRPS—vary significantly.
-This suggests that confidence evaluation must be treated as a separate challenge from prediction assessment.
-Boolean Questions:
-Table 4 shows that while some models achieve reasonable accuracy, their calibration remains inconsistent.
-For example, GPT2-XL achieves an accuracy of 0.67 but has a relatively high Brier score of 0.45, while OLMo-7B-Instruct achieves similar accuracy (0.65) with a lower Brier score of 0.38, suggesting better confidence estimation.
-Notably, models with later training cutoffs do not always outperform older ones; for instance, LLaMA-7B has lower accuracy (0.57) and a worse Brier score (0.52) than some earlier models.
-This indicates that forecasting accuracy depends on multiple factors beyond knowledge recency.
-Model N ADE ($\downarrow$) CRPS ($\downarrow$)
-GPT2 26 0.99 0.99
-GPT2-XL 26 1.00 1.00
-Pythia-14m 25 1.00 1.00
-Pythia-160m 25 1.00 1.00
-Pythia-2.8b 25 0.97 0.96
-Bloom-560m 12 1.00 1.00
-Bloom-7b1 12 1.00 1.00
-Llama-7b 10 0.98 0.98
-OLMo-1B 6 1.00 1.00
-OLMo-7B 6 1.00 1.00
-OLMo-7B-Instruct 6 1.00 1.00
-OLMo-2-7B 4 1.00 1.00
-OLMo-2-7B-Instruct 4 0.87 0.82
-Table 5: Performance of forecasting models on Timeframe Prediction tasks in FOReCAst.
-Metrics include the number of evaluated questions (N), the normalized absolute days error (ADE), and the Continuous Ranked Probability Score (CRPS), where lower values indicate better performance.
-Model N APE ($\downarrow$) MAE ($\downarrow$) CRPS ($\downarrow$)
-GPT2 81 0.23 0.87 0.86
-GPT2-XL 81 0.03 0.75 0.72
-Pythia-14m 77 0.17 0.88 0.85
-Pythia-160m 77 0.03 0.78 0.76
-Pythia-2.8b 77 0.07 0.81 0.79
-Bloom-560m 43 0.05 0.73 0.71
-Bloom-7b1 43 0.06 0.75 0.72
-Llama-7b 32 0.06 0.64 0.61
-OLMo-1B 23 0.21 0.79 0.76
-OLMo-7B 23 0.22 0.84 0.82
-OLMo-7B-Instruct 23 0.24 0.80 0.77
-OLMo-2-7B 20 0.02 0.65 0.62
-OLMo-2-7B-Instruct 20 0.09 0.60 0.57
-Table 6: Performance of forecasting models on Quantity Estimation tasks in FOReCAst.
-Metrics include the number of evaluated questions (N), normalized absolute percentage error (APE), mean absolute error (MAE), and Continuous Ranked Probability Score (CRPS), where lower values are better.
-Timeframe Prediction:
-Table 5 reveals that most models struggle with predicting event timing, as both ADE and CRPS remain near the worst-case scenario of 1.0.
-Across all models, only one achieves a CRPS below 0.9: OLMo-2-7B-Instruct (CRPS = 0.82), which still indicates substantial uncertainty.
-Even models with more recent training data, such as OLMo-2-7B (CRPS = 0.99), fail to make well-calibrated temporal forecasts.
-These results suggest that even when models correctly anticipate whether an event will happen, quantifying when it will occur remains a major challenge.
-Quantity Estimation:
-Table 6 highlights that models differ widely in their ability to estimate numerical values.
-Some achieve relatively low APE, such as GPT2-XL (APE = 0.03) and Pythia-160M (APE = 0.02), yet their confidence calibration, CRPS, does not always align with their point prediction performance.
-For instance, GPT2-XL has a CRPS of 0.72, while Pythia-160M has a slightly higher CRPS of 0.76, despite achieving lower APE.
-In contrast, OLMo-2-7B-Instruct, which has a relatively higher APE of 0.08, achieves the lowest CRPS (0.57) among all models.
-These results indicate that prediction quality and confidence calibration do not necessarily improve together, reinforcing the complexity of numerical forecasting.
+5.2 Results, Findings, and Discussions
+Our experiments on the FOReCAst dataset reveal that forecasting remains a challenging task for current LLMs, particularly in estimating and expressing uncertainty.
+While models often achieve reasonable accuracy in their point predictions, their confidence calibration—measured by Brier score and CRPS—shows substantial variability.
+This indicates that uncertainty estimation should be treated as a distinct evaluation axis, not inferred from accuracy alone.
+Throughout the following sections, we offer interpretations and possible explanations for the observed behaviors.
+Boolean Questions
+Table 4 shows that while some models reach reasonable accuracy on binary classification tasks, their confidence calibration remains inconsistent.
+For instance, GPT2-XL achieves a relatively strong accuracy of 0.67 but has a higher Brier score of 0.45, whereas OLMo-7B-Instruct achieves a similar accuracy (0.65) with a lower Brier score of 0.38, indicating better alignment between predicted probabilities and actual outcomes.
+This discrepancy suggests that models may differ not only in their ability to identify the correct answer, but also in how they distribute confidence over binary options.
+Importantly, model performance does not appear to correlate straightforwardly with scale or training recency.
+For example, LLaMA-7B, a more recent model, shows lower accuracy (0.57) and higher Brier score (0.52) than older models like GPT2-XL.
+This supports the observation that accuracy in binary forecasting tasks depends on a combination of world knowledge, probabilistic reasoning, and how well models can handle epistemic uncertainty, particularly when questions are unresolved or partially underspecified at prompt time.
+The gap between classification accuracy and calibration further underscores the importance of evaluating both, as good classification alone may mask unreliable probability estimates.
+Quantity Estimation Timeframe Prediction
+Model N APE ($\downarrow$) MAE ($\downarrow$) CRPS ($\downarrow$) Model N ADE ($\downarrow$) CRPS ($\downarrow$)
+GPT2 81 0.23 0.87 0.86 GPT2 26 0.99 0.99
+GPT2-XL 81 0.03 0.75 0.72 GPT2-XL 26 1.00 1.00
+Pythia-14m 77 0.17 0.88 0.85 Pythia-14m 25 1.00 1.00
+Pythia-160m 77 0.03 0.78 0.76 Pythia-160m 25 1.00 1.00
+Pythia-2.8b 77 0.07 0.81 0.79 Pythia-2.8b 25 0.97 0.96
+Bloom-560m 43 0.05 0.73 0.71 Bloom-560m 12 1.00 1.00
+Bloom-7b1 43 0.06 0.75 0.72 Bloom-7b1 12 1.00 1.00
+Llama-7b 32 0.06 0.64 0.61 Llama-7b 10 0.98 0.98
+OLMo-1B 23 0.21 0.79 0.76 OLMo-1B 6 1.00 1.00
+OLMo-7B 23 0.22 0.84 0.82 OLMo-7B 6 1.00 1.00
+OLMo-7B-Instruct 23 0.24 0.80 0.77 OLMo-7B-Instruct 6 1.00 1.00
+OLMo-2-7B 20 0.02 0.65 0.62 OLMo-2-7B 4 1.00 1.00
+OLMo-2-7B-Instruct 20 0.09 0.60 0.57 OLMo-2-7B-Instruct 4 0.87 0.82
+Table 5: Performance of forecasting models on Quantity Estimation (left) and Timeframe Prediction (right) tasks in FOReCAst.
+All metrics are lower-is-better.
+APE: Absolute Percentage Error, MAE: Mean Absolute Error, CRPS: Continuous Ranked Probability Score, ADE: Absolute Days Error.
+Timeframe Prediction
+Table 5 demonstrates that timeframe prediction—estimating when an event will occur—is especially difficult for current models.
+Across the board, both ADE and CRPS values remain close to 1.0, indicating little improvement over naive or uninformed baselines.
+Only OLMo-2-7B-Instruct achieves a CRPS noticeably below this range (0.82), but even this is far from well-calibrated.
+The generally poor performance suggests that temporal forecasting presents challenges beyond simple factual retrieval or classification.
+This difficulty likely reflects the inherent complexity of temporal prediction, which often requires models to reason over event dynamics, causality, and elapsed time—all aspects that are underrepresented in standard pretraining corpora.
+In contrast to Boolean tasks, where resolution criteria are clearly defined and outcomes are binary, timeframe questions require mapping language input to a continuous temporal target, which demands a finer-grained understanding of how events evolve over time.
+That most models fail to depart meaningfully from baseline-level CRPS indicates that LLMs may lack sufficient temporal inductive biases or exposure to structured time-referenced reasoning.
+Quantity Estimation
+As shown in Table 5, quantity estimation—predicting a numerical outcome—exhibits a broader range of performance across models.
+Some models such as Pythia-160M and GPT2-XL achieve very low absolute percentage error (APE = 0.02 and 0.03 respectively), indicating strong point prediction capability.
+However, their CRPS values (0.76 and 0.72) suggest relatively poor calibration.
+In contrast, OLMo-2-7B-Instruct, with a higher APE of 0.08, attains the lowest CRPS (0.57), suggesting it better captures uncertainty even if its point estimates are less precise.
+This divergence illustrates that point prediction accuracy and confidence calibration are not necessarily aligned.
+Predicting a correct numerical value does not guarantee that a model’s probability distribution around its prediction reflects uncertainty appropriately.
+Since CRPS penalizes both distance from the true value and over/under-confidence, this metric provides a more comprehensive view of performance in numerical forecasting tasks.
+These results suggest that even relatively strong predictive models may require additional mechanisms—such as structured uncertainty modeling or calibration-sensitive objectives—to produce well-calibrated forecasts in quantity estimation settings.
 Impact of Model Size on Forecasting Performance
-Larger models do not consistently improve forecasting performance.
-Within the Pythia family, Pythia-2.8b shows occasional gains in accuracy and calibration over Pythia-14m and Pythia-160m, but the improvements are not uniform across all metrics.
-Similarly, while OLMo-2-7B variants sometimes achieve lower CRPS and MAE in quantity estimation, these benefits often come with trade-offs in point prediction.
-These results suggest that model size alone is not a reliable predictor of forecasting performance.
-This highlights the importance of developing task-specific techniques rather than relying on ever-larger models to solve the forecasting problem.
+Model size alone does not reliably predict forecasting performance across tasks.
+Within the Pythia family, Pythia-2.8b occasionally improves over smaller variants in certain metrics—for example, it achieves lower CRPS than Pythia-14m on quantity estimation—but these gains are neither large nor consistent.
+In some cases, smaller models like Pythia-160M outperform their larger counterparts in point prediction metrics such as APE.
+A similar trend appears within the OLMo family: while the OLMo-2-7B variants perform competitively in calibration metrics (e.g., CRPS), this is not accompanied by uniform improvements in accuracy or absolute error.
+These observations suggest that increases in parameter count do not automatically translate to improved forecasting performance.
+While larger models often benefit from greater representational capacity, the forecasting task—particularly when it involves expressing well-calibrated uncertainty—may place demands that are orthogonal to model size.
+Specifically, forecasting tasks may require targeted reasoning over time, quantities, or epistemic uncertainty, which are not clearly emphasized during standard pretraining or scaled parameter growth.
+These findings indicate that effective forecasting may depend more on model behavior during decoding or task-specific adaptation, rather than scale alone.
 Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
@@
 For quantity estimation, instruct-tuned models have slightly higher APE but lower MAE and CRPS, suggesting better uncertainty calibration.
-These results indicate that instruction tuning enhances confidence estimation, even if it does not always improve point prediction accuracy.
-This suggests a trade-off where instruct-tuned models prioritize more reliable uncertainty quantification.
+While these results point to potential benefits of instruction tuning for probabilistic reasoning, the sample size remains limited due to the recency of instruction-tuned models, and the observed gains may not be statistically robust.
 Aggregation Method N Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$)
@@
 Bayesian Aggregation achieves the highest accuracy (0.58) and F1 score (0.35), suggesting it is the most effective at identifying correct outcomes.
-However, Weighted Average yields a significantly lower Brier score (0.2914), indicating superior confidence calibration compared to other methods.
+However, Weighted Average yields a significantly lower Brier score (0.29), indicating superior confidence calibration compared to others.
 Majority Vote, Highest Confidence, and Logit Mean Probability produce comparable accuracy and F1 scores but have noticeably higher Brier scores, suggesting weaker uncertainty estimation.
@@
 The challenge remains in developing techniques that optimize both accuracy and calibration simultaneously, emphasizing the importance of uncertainty-aware forecasting.
+In preliminary experiments, we also explored verbalized confidence, prompting the model to explicitly generate a probability (e.g., “I’m 80% confident…”).
+However, most models either failed to produce enough meaningful numeric estimates or generated inconsistent responses, particularly on harder questions.
+This highlights a limitation in current instruction-following capabilities when it comes to quantifying uncertainty directly.
+A promising direction for future work is to fine-tune models specifically to produce structured confidence outputs, possibly by supervising against calibrated targets or training with ranking-based objectives.
+Additionally, integrating task-specific priors (e.g., event resolution frequency, temporal proximity) into aggregation could further improve calibration.
+Hence, joint modeling of correctness and uncertainty remains a key open challenge for real-world deployment of forecasting systems.
+Lack of Correlation Between Forecast Date and Model Performance
+One might expect language models to perform better on forecasting questions that are closer to their training cutoff date, since more relevant information may have been available during pretraining.
+For instance, a model like GPT2, with a cutoff in December 2017, would intuitively have an advantage when forecasting 2018 events compared to events in 2020.
+However, as shown in Appendix E, we observe no clear improvement for near-cutoff questions.
+In many cases, model performance remains flat or even inconsistent across time, with no systematic degradation as the temporal distance increases.
+This observation suggests that forecasting is meaningfully different from retrieval-based tasks such as fact-checking or question answering.
+While those tasks often benefit from training data that includes the target facts or related content, forecasting involves reasoning under uncertainty about events that had not yet occurred during training.
+Models are not explicitly trained to perform such reasoning, and pretraining objectives generally do not reward predictions about future developments.
+The absence of a correlation between cutoff proximity and performance highlights that forecasting requires capabilities beyond knowledge access, including abstraction, probabilistic reasoning, and temporal inference.
+These factors make the task more complex and reinforce the need for dedicated methods that go beyond static knowledge retrieval.
 6 Related Work
 Recent forecasting benchmarks focus on event prediction but largely overlook confidence calibration.
-OpenForecast (Wang et al., 2025) introduces a large-scale dataset for open-ended, multi-step event forecasting but does not assess model confidence.
-ForecastBench (Karger et al., 2024) evaluates binary (Yes/No) forecasting by prompting models for direct probability estimates, but since it queries each option separately, the assigned probabilities do not necessarily sum to 1, leading to potential inconsistencies.
-Neither benchmark systematically evaluates confidence calibration, a crucial aspect for reliable forecasting in real-world applications.
+OpenForecast (Wang et al., 2025) introduces a large-scale dataset for open-ended, multi-step event forecasting but does not assess model confidence or calibration.
+ForecastBench (Karger et al., 2024) evaluates binary (Yes/No) forecasting by prompting models for direct probability estimates, but since it queries each option independently, the assigned probabilities do not necessarily sum to 1, leading to potential inconsistencies.
+Both benchmarks advance the study of future event prediction but stop short of systematically evaluating uncertainty quantification, a crucial aspect for reliable deployment in decision-making contexts.
+Our work complements these efforts by providing a structured setting for evaluating both forecasting accuracy and probabilistic calibration.
 Beyond forecasting, several benchmarks assess language models’ reasoning and inference capabilities.
 COPA (Roemmele et al., 2011) evaluates causal reasoning by presenting a premise and two alternatives, requiring models to select the more plausible cause or effect.
-HellaSwag Zellers et al.
-(2019) challenges models with sentence completion tasks that demand commonsense reasoning, where models must choose the most sensible continuation of a given scenario.
+HellaSwag (Zellers et al., 2019) challenges models with sentence completion tasks that demand commonsense reasoning, where models must choose the most sensible continuation of a given scenario.
 PRobELM (Yuan et al., 2024b) assesses models’ capacity to rank scenarios by plausibility, bridging the gap between factual accuracy and world knowledge.
-While these benchmarks provide insights into models’ reasoning abilities, they do not address the challenges of forecasting future events.
+While these benchmarks probe useful reasoning skills, they are not designed for evaluating predictions about unresolved or future outcomes, nor do they assess confidence calibration in a dynamic context.
+7 Limitations
+FOReCAst is constructed entirely from Metaculus, a platform known for its high-quality, well-resolved forecasting questions and active user base.
+This makes it a strong and consistent foundation for developing the task of probabilistic forecasting with LLMs.
+However, it does not cover the full diversity of forecasting practices, such as different question formats, domain emphases, or forecasting styles seen in other platforms like Good Judgment Open or domain-specific expert settings.
+In this work, our primary focus is to define the task rigorously and establish a benchmark that is methodologically clean and scalable.
+Expanding to additional data sources and platforms is part of our future work and should help test generalization across forecasting cultures and environments.
+Our method for constructing gold confidence labels is based on community-aggregated forecasts, which offer a practical and scalable proxy for crowd belief and uncertainty.
+While this approach enables systematic evaluation across hundreds of questions, it may not fully capture the variability and richness of individual forecasters’ reasoning or levels of disagreement.
+Moreover, the aggregated predictions may sometimes be influenced by group dynamics or correlated sources of error.
+We acknowledge these trade-offs, and view our current approach as a reasonable starting point given the limited availability of high-quality, large-scale confidence annotations.
+Future work could explore incorporating richer uncertainty representations, such as forecast distributions, forecaster rationales, or disagreement metrics.
+Finally, our dataset is currently limited to English-language questions and forecasts, reflecting the language used by Metaculus and many existing NLP resources.
+While this is a reasonable initial scope for establishing the task, we recognize that forecasting behavior, uncertainty expression, and interpretation of probabilities may vary significantly across linguistic and cultural contexts.
+Extending FOReCAst to include multilingual questions and community forecasts is a promising direction for improving the inclusiveness and robustness of future benchmarks.
+We also note that some dimensions, such as temporal adaptation and real-time reasoning, remain out of scope here but are essential long-term goals for building forecasting-capable systems.
 8 Conclusion
@@
 Our results show that current models struggle with both prediction and well-calibrated confidence, underscoring the need for improved uncertainty estimation and confidence calibration.
-Limitations
-While FOReCAst represents a significant step toward evaluating forecasting accuracy and confidence calibration, there are inherent limitations that we view as opportunities for future research rather than fundamental shortcomings of our work.
-First, our dataset is constructed solely from Metaculus, which may not fully represent the global diversity of forecasting practices or question domains.
-Second, our method for deriving gold confidence relies on community forecasts and heuristic transformations that might not capture all nuances of human uncertainty.
-Lastly, our focus on English-language forecasts limits the benchmark’s applicability across different languages and cultural contexts.
-Addressing these issues is part of our future work agenda.
-Ethical Statement
-FOReCAst is built from data sourced exclusively from Metaculus, an English-language forecasting platform.
-As a result, the dataset may embody the linguistic, cultural, and social biases inherent in its user community.
-These biases could affect both question selection and confidence judgments.
-We acknowledge these concerns and stress that our benchmark is intended as an initial step toward more inclusive forecasting evaluations.
-Future efforts should aim to incorporate data from a broader range of platforms and languages to mitigate these biases.
+Our dataset is now publicly available.
 Acknowledgements
@@
 Table 8: Dataset statistics for FOReCAst, showing the distribution of questions across different forecasting types, with the overall total in the last column.
+Table 9: Top 10 months and categories across the full dataset and each question type. Note that questions can belong to multiple categories, reflecting the diversity of domains represented.
+Global Distribution Timeframe Prediction
+Month Count % Month Count %
+2023-12 187 8.29% 2021-05 9 6.47%
+2024-12 105 4.65% 2021-03 7 5.04%
+2022-12 89 3.95% 2022-12 6 4.32%
+2021-01 65 2.88% 2024-01 6 4.32%
+2024-01 62 2.75% 2022-02 6 4.32%
+2024-05 56 2.48% 2021-04 5 3.60%
+2025-01 49 2.17% 2024-05 4 2.88%
+2023-01 48 2.13% 2022-09 4 2.88%
+2020-11 47 2.08% 2021-12 4 2.88%
+2020-03 41 1.82% 2021-01 4 2.88%
+Category Count % Category Count %
+Politics 571 17.03% Health & Pandemics 40 18.10%
+Geopolitics 486 14.50% Economy & Business 31 14.03%
+Economy & Business 476 14.20% Politics 26 11.76%
+Health & Pandemics 313 9.34% Technology 23 10.41%
+Technology 253 7.55% Computing and Math 21 9.50%
+Elections 218 6.50% Geopolitics 19 8.60%
+Computing and Math 192 5.73% Artificial Intelligence 17 7.69%
+Artificial Intelligence 151 4.50% Space 13 5.88%
+Natural Sciences 139 4.15% Law 7 3.17%
+Law 104 3.10% Natural Sciences 6 2.71%
+Quantity Estimation Boolean Questions
+Month Count % Month Count %
+2021-02 25 6.96% 2023-12 167 9.50%
+2022-12 21 5.85% 2024-12 94 5.35%
+2020-04 19 5.29% 2022-12 62 3.53%
+2023-12 18 5.01% 2024-01 51 2.90%
+2024-05 17 4.74% 2021-01 45 2.56%
+2020-03 17 4.74% 2025-01 44 2.50%
+2021-01 16 4.46% 2023-01 38 2.16%
+2020-11 16 4.46% 2024-05 35 1.99%
+2021-10 12 3.34% 2023-06 34 1.93%
+2020-12 12 3.34% 2024-11 29 1.65%
+Category Count % Category Count %
+Economy & Business 102 22.57% Politics 488 18.22%
+Health & Pandemics 88 19.47% Geopolitics 428 15.98%
+Politics 57 12.61% Economy & Business 343 12.80%
+Geopolitics 39 8.63% Technology 205 7.65%
+Artificial Intelligence 27 5.97% Elections 190 7.09%
+Elections 25 5.53% Health & Pandemics 185 6.91%
+Technology 25 5.53% Computing and Math 151 5.64%
+Computing and Math 20 4.42% Natural Sciences 124 4.63%
+Environment & Climate 16 3.54% Artificial Intelligence 107 3.99%
+Sports & Entertainment 13 2.88% Law 96 3.58%
 Table 8 presents detailed dataset statistics, including the total number of questions and their distribution across Boolean Questions, Timeframe Prediction, and Quantity Estimation tasks.
+Table 9 provides the top 10 categories and closing months for each task type, illustrating the breadth of topics and temporal diversity in FOReCAst.
 Appendix C Prompts
@@
 D.2 Evaluation Hyperparameters
-The scaling factor $\alpha$ in Equation 4, Equation 8, and Equation 9 is set to 0.05.
-For Equation 5 and Equation 6, we set $\sigma_{\max}$ to 30 and $\sigma_{\min}$ to 1 for Timeframe Prediction, while for Quantity Estimation, $\sigma_{\max}$ is 20 and $\sigma_{\min}$ is 1.
+The scaling factor $\alpha$ in Equation 3 and Equation 6 is set to 0.05.
+For Equation 4, we set $\sigma_{\max}$ to 30 and $\sigma_{\min}$ to 1 for Timeframe Prediction, while for Quantity Estimation, $\sigma_{\max}$ is 20 and $\sigma_{\min}$ is 1.
 These values ensure that evaluation metrics appropriately scale errors and confidence calibration.

### v4

@@
 Abstract
-Forecasting is an important task in many domains, such as technology and economics.
-However existing forecasting benchmarks largely lack comprehensive confidence assessment, focus on limited question types, and often consist of artificial questions that do not align with real-world human forecasting needs.
-To address these gaps, we introduce FOReCAst (Future Outcome Reasoning and Confidence Assessment), [^1] Publicly availble at https://huggingface.co/datasets/MoyYuan/FOReCAst a benchmark that evaluates models’ ability to make predictions and their confidence in them.
+Forecasting is an important task in many domains.
+However, existing forecasting benchmarks lack comprehensive confidence assessment, focusing on limited question types, and often consist of artificial questions that do not reflect real-world needs.
+To address these gaps, we introduce FOReCAst (Future Outcome Reasoning and Confidence Assessment), a benchmark that evaluates models’ ability to make predictions and their confidence in them.
 FOReCAst spans diverse forecasting scenarios involving Boolean questions, timeframe prediction, and quantity estimation, enabling a comprehensive evaluation of both prediction accuracy and confidence calibration for real-world applications.
+[^1] FOReCAst is publicly available at https://huggingface.co/datasets/MoyYuan/FOReCAst.
 1 Introduction
 Recent advances in large language models (LLMs) have significantly improved their performance across a wide range of natural language processing (NLP) tasks.
-Alongside these developments, various benchmarks and datasets have been introduced to effectively assess the capabilities of LLMs, particularly in terms of knowledge and reasoning (Zellers et al., 2019; Guo et al., 2023).
-Fact-based benchmarks, such as TruthfulQA (Lin et al., 2022), evaluate LLMs based on factual correctness, focusing on tasks like retrieving and verifying facts that are already known.
+In parallel to these developments, various benchmarks and datasets have been introduced to effectively assess the capabilities of LLMs, particularly in terms of knowledge and reasoning (Zellers et al., 2019; Guo et al., 2023).
+Fact-oriented benchmarks, such as TruthfulQA (Lin et al., 2022), evaluate LLMs on factual correctness, focusing on tasks like retrieving and verifying facts that are already known.
 Forecasting is a crucial yet challenging task across various domains, including technology, economics, and public policy.
 Unlike tasks that rely on retrieving and verifying existing knowledge, forecasting requires predicting plausible outcomes for future events, often under uncertainty and incomplete information.
-This makes forecasting particularly difficult, as models must infer trends, assess probabilities, and adapt to new information.
 Several datasets have been introduced to evaluate LLMs’ forecasting capabilities.
 ForecastQA (Jin et al., 2020) uses a multiple-choice format where models predict future outcomes, but it lacks confidence assessment.
-AutoCast (Zou et al., 2022) incorporates confidence intervals; however, its confidence estimates are not specifically designed for forecasting tasks.
-Other datasets, such as ExpTime (Yuan et al., 2024a), are artificially generated from structured data and focus on explainable event forecasting based on temporal knowledge graphs.
+AutoCast (Zou et al., 2022) incorporates confidence intervals but not for its forecasting questions.
+Other datasets, such as ExpTime (Yuan et al., 2024a), are artificially generated from temporal knowledge graph, focusing on explainability.
 All of these aforementioned benchmarks overlook a crucial aspect of forecasting: confidence evaluation.
-Confidence plays a central role in forecasting, as predictions about unresolved events inherently lack definitive correctness at the time of evaluation.
+Confidence plays a central role in forecasting, as predictions about unresolved events inherently lack definitive answers at the time they are asked.
 Predictions made with absolute certainty are undesirable, even if they ultimately prove to be correct, because they fail to account for the uncertain nature of future events.
@@
 Therefore, well-calibrated confidence scores are as crucial as the accuracy of the predictions themselves.
-To address these gaps, we present FOReCAst: Future Outcome Reasoning and Confidence Assessment.
-FOReCAst focuses on three distinct types of forecasting questions, shown in Table 1: (1) Boolean Questions, such as "Will there be a Frontier AI lab in China before 2026?"; (2) Timeframe Prediction, such as "When will OpenAI announce GPT-5?"; and (3) Quantity Estimation, such as "How many spacecrafts will land on the moon in each of the following years?" We conduct experiments using a range of models differing in size, training objectives, and training cutoff times, and explore multiple methods for estimating model confidence.
-Our results reveal that forecasting remains highly challenging for current LLMs, particularly in confidence evaluation, with no consistent correlation between prediction accuracy and calibration.
+To address these issues we present FOReCAst: Future Outcome Reasoning and Confidence Assessment.
+Gold confidence scores are derived from aggregated human forecasts, reflecting community consensus over time.
+FOReCAst focuses on three distinct types of forecasting questions, as illustrated in Table 1: (1) Boolean Questions, such as "Will there be a Frontier AI lab in China before 2026?"; (2) Timeframe Prediction, such as "When will OpenAI announce GPT-5?"; and (3) Quantity Estimation, such as "How many spacecrafts will land on the moon in 2025?" We conduct experiments using a range of models differing in size, training objectives, and training cutoff times, and explore multiple methods for estimating model confidence.
+Our results reveal that forecasting remains highly challenging for contemporary LLMs.
+Their confidence calibration is poor overll, and it has inconsistent correlation with predictive accuracy.
 While larger models sometimes improve performance, the effect is inconsistent and highly task-dependent.
+Table 1: Forecasting questions with their resolutions and confidence scores.
 Type Question Resolution Confidence
-Boolean Question Will a Frontier AI lab be established in China by 2026? Yes 0.73
+Boolean Question Will there be a Frontier AI lab in China before 2026? Yes 0.73
 Timeframe Prediction When will OpenAI announce GPT-5? 2024-08-01 0.85
 Quantity Estimation How many spacecrafts will land on the moon in 2025? 3 0.65
-Table 1: Examples of forecasting questions with their resolutions and confidence scores.
 2 FOReCAst: Problem Formulation
-System responses in FOReCAst consist of (1) a prediction answering a question given the available information and (2) a confidence score in the prediction.
+In FOReCAst a system response consists of (1) a prediction that answers a question given the available information and (2) a confidence score in the prediction.
 This ensures a comprehensive assessment of forecasting, accounting for both correctness and confidence calibration.
@@
 (1) Boolean Questions, which ask yes/no questions about the occurrence of future events (sometimes within a certain timeframe).
-Boolean questions are simple to evaluate, and they can still be surprisingly challenging  (Clark et al., 2019).
+Even if such questions are simple to evaluate, they can still be surprisingly challenging  (Clark et al., 2019).
 (2) Timeframe Prediction, which requires predicting a specific timeframe for an event, and are essential for applications where knowing whether an event will happen or not without a timeframe is insufficient.
-(3) Quantity Estimation, which involves providing numerical estimates related to future events, e.g., economic indicators or trends.
+(3) Quantity Estimation, which involves providing numerical estimates related to future events.
 Formally, let $Q$ represent a question about a future event, and let $M$ denote a system with access to information up to time $t_{train}$ (e.g., the system’s knowledge cutoff point).
@@
 3 Evaluating Predictions and Confidence
-Evaluating Boolean Questions
+Boolean Questions
 For Boolean questions, where the answer space is $\mathcal{A}=\{\text{Yes},\text{No}\}$, prediction performance is evaluated using standard classification metrics, including accuracy and F1-score.
@@
 where $C^{\text{pred}}_{i}$ is the model’s predicted confidence and $C^{\text{gold}}_{i}$ is the gold confidence.
-This modification ensures that models are evaluated based on their ability to match the likelihood of an event.
-Lower Brier scores indicate better calibration, reflecting how well the predicted confidence aligns with the likelihood of the event.
-Evaluating Timeframe Prediction
+This modification ensures that models are evaluated based on their ability to match the likelihood of an event, with lower Brier scores indicating better calibration.
+Timeframe Prediction
 For timeframe prediction, where the answer space consists of specific dates in the YYYY-MM-DD format, predictive accuracy is measured using absolute day error (ADE).
@@
 where $\alpha$ is a scaling factor that controls how sharply large errors are penalized. This transformation ensures that extreme deviations do not disproportionately dominate the evaluation.
-For confidence calibration, rely on the Continuous Ranked Probability Score (CRPS) (Matheson and Winkler, 1976), which is a generalisation of the Mean Absolute Error to probabilistic forecasts, and extend it to compare the predicted probability distribution with a gold distribution.
-Specifically, we assume that both the predicted and the gold confidence predictions follow Gaussian distributions, namely $\mathcal{N}(D^{\text{pred}}_{i},\sigma^{\text{pred}}_{i})$ and $\mathcal{N}(D^{\text{gold}}_{i},\sigma^{\text{gold}}_{i})$ respectively, where the standard deviations are computed as:
+For confidence calibration, we rely on the Continuous Ranked Probability Score (CRPS) (Matheson and Winkler, 1976), which is a generalisation of the Mean Absolute Error to probabilistic forecasts, and extend it to compare the predicted probability distribution with a gold distribution.
+Specifically, we assume that both the predicted and the gold confidence predictions follow Gaussian distributions $\mathcal{N}(D^{\text{pred}}_{i},\sigma^{\text{pred}}_{i})$ and $\mathcal{N}(D^{\text{gold}}_{i},\sigma^{\text{gold}}_{i})$ respectively, where the standard deviations are computed as:
 $\sigma^{\text{pred}}_{i}=\sigma_{\max}\cdot(1-C^{\text{pred}}_{i})+\sigma_{\min}\cdot C^{\text{pred}}_{i},\quad\sigma^{\text{gold}}_{i}=\sigma_{\max}\cdot(1-C^{\text{gold}}_{i})+\sigma_{\min}\cdot C^{\text{gold}}_{i}.$ (4)
@@
 A lower CRPS indicates better calibration, as it reflects a closer match between the predicted uncertainty and the uncertainty as specified by the gold confidence.
-Evaluating Quantity Estimation
+Quantity Estimation
 For quantity estimation, where the answer space consists of non-negative real numbers ($\mathcal{A}=\mathbb{R}_{\geq 0}$), we evaluate prediction performance using two error metrics: absolute percentage error (APE) and mean absolute error (MAE).
@@
 4.1 Data Source and Question Selection
-FOReCAst is constructed from Metaculus,222www.metaculus.com.
-Examples in Appendix A.
-an online forecasting platform where forecasters submit probabilistic predictions to questions across various domains.
-Metaculus aggregates individual forecasts into a continuously updated community prediction, which is finalized just before resolution.
-Each question has predefined resolution criteria, ensuring verifiable outcomes.
-To ensure dataset reliability, we include only questions with a definitive resolution and at least 100 forecasts to maintain statistical reliability.
-Ambiguous or subjectively resolved questions are excluded, and we remove those whose outcomes depend on arbitrary or uncontrollable factors.
-These steps ensure that FOReCAst consists of high-quality, well-formed forecasting questions with verifiable outcomes.
-The dataset spans a diverse range of domains such as politics, economics, health, and technology, with questions distributed across a wide temporal range; detailed statistics are provided in Table 9.
+FOReCAst is constructed from Metaculus (www.metaculus.com, see examples in Appendix A), an online forecasting platform where users submit forecasts to questions across domains such as politics, economics, health, and technology.
+Metaculus aggregates these into a community prediction that is continuously updated until shortly before resolution.
+Each question has predefined resolution criteria, and the platform enforces guidelines to ensure that all predictions comply with them and are human-generated.
+We discard questions that Metaculus marks as ambiguous or those resolved in ways that make forecasting infeasible—for example, when the outcome falls outside the prediction range or when resolution criteria change after submission.
+Also, we only include questions with at least 100 forecasts to maintain statistical reliability.
 4.2 Extracting Confidence from Crowdsourced Forecasts
-Forecasting distinguishes between physical probabilities—objective likelihoods derived from statistical or scientific models—and human beliefs (Sanders, 1963) about future events.
-While physical probabilities can be useful, they are often unavailable, particularly for questions involving human behavior, economics, or sociopolitical outcomes.
-Instead, collective human forecasts offer a more practical confidence estimate, integrating expert reasoning, contextual knowledge, and evolving evidence.
-For instance, predicting a technological breakthrough depends more on expert assessment and current trends than on rigid probabilistic models.
-Therefore, confidence in FOReCAst is derived from Metaculus community forecasts, which aggregate predictions from a diverse pool of forecasters.
-While human predictions are sometimes incorrect, they still serve as a valuable proxy for uncertainty, as they reflect the best available reasoning given the information at the time.
-Formal Definition of Gold Confidence
+In an ideal scenario, we would have both the resolutions and associated confidence scores available at the time a forecast question is posed.
+However, in practice, we only observe questions with their eventual resolutions and crowd-sourced human forecasts.
+To derive confidence scores we aggregate these human forecasts, which, despite occasional disagreements especially on non-trivial questions, provide informative signals about confidence.
+Human predictions can be sometimes incorrect, but they can still serve as a valuable proxy for uncertainty, as they reflect the best available reasoning given the information at the time.
+Metaculus ensures that these forecasts are manually submitted and fall within calibrated ranges, which encourages us to use them as proxies for evaluating model confidence.
+This is inline with other approaches to benchmark construction on tasks where humans disagree on labels, e.g.
+ChaosNLI (Nie et al., 2020; Plank, 2022).
 Gold confidence in FOReCAst is derived from the final Metaculus community prediction just before question resolution.
-Rather than directly using the predicted probability for the correct outcome, we compute a log score relative to a uniform baseline, following Metaculus’s established approach for evaluating probabilistic forecasts  [^3] https://www.metaculus.com/help/scores-faq/.
+Following Metaculus’s established approach for evaluating probabilistic forecasts,  [^2] https://www.metaculus.com/help/scores-faq/ we compute a log score relative to a uniform baseline.
 This transformation ensures that confidence reflects not only correctness, but also the amount of informational gain over random guessing.
@@
 This approach accounts for the resolution granularity of each question and ensures confidence is not inflated in uncertain or diffuse distributions.
-While this scoring formulation is tailored to Metaculus, it can be adapted or replaced if used with other forecasting platforms or scoring protocols.
-Our goal is to align confidence with meaningful deviations from uncertainty, while maintaining a consistent and interpretable signal across all tasks.
 4.3 Dataset Statistics and Comparison
@@
 Each question includes a resolved outcome, a gold confidence score, and a final Metaculus community forecast before resolution.
+The questions cover a broad temporal range, with resolutions spanning from March 2016 to the end of January 2025.
 To facilitate model development and evaluation, we split the dataset into 65% training, 10% validation, and 25% test.
 The full dataset statistics is shown in Appendix B.
+Note that we do not use the training split for model fine-tuning or prompt optimization; all models are evaluated in a one-shot setting.
 Table 2 provides a comparison between FOReCAst and existing forecasting benchmarks.
 Compared to prior datasets, FOReCAst uniquely emphasizes both forecasting accuracy and confidence calibration, includes a diverse set of forecasting tasks, and is constructed from a well-established crowdsourced platform with rigorous resolution criteria.
+Table 2: Comparison of key features across our benchmark variants, highlighting our evaluation of confidence across different question types.
 Benchmark Question Types Natural Questions Confidence
@@
 FOReCAst Various
-Table 2: Comparison of key features across our benchmark variants, highlighting our evaluation of confidence across different question types.
+Extensibility and Long-Term Sustainability
+FOReCAst is built with extensibility and sustainability in mind.
+Our data collection pipeline is fully automated, enabling continuous updates as new forecasting questions and outcomes become available.
+Although this release draws from Metaculus due to its high-quality resolution criteria and historical depth, the same pipeline can be adapted to other platforms such as Polymarket, Manifold Markets, or Good Judgment Open.
+By supporting multiple data sources and maintaining alignment with evolving real-world events, FOReCAst provides a sustainable foundation for benchmarking forecasting models over time.
+This extensibility not only ensures long-term utility, but also enables future work to study model generalization across domains, platforms, and temporal shifts.
 5 Experiments on FOReCAst
-5.1 Experimental Setup
-Models.
+5.1 Models
 We evaluate a diverse set of large language models (LLMs) with varying training data cutoffs, model sizes, and instruction tuning.
-To analyze the impact of knowledge recency, we group models by family and assume the cutoff date is the 1st of the stated month.
-The models used are shown in Table 3.
+To analyze the impact of knowledge recency, we group models by family and use the stated month in Table 3 as the assumed cutoff.
+[^3] The exact cutoff date of Qwen 2.5 is unclear, with sources suggesting dates ranging from October 2023 to December 2023; we conservatively use 2023-12 as the assumed cutoff to avoid potential knowledge leakage.
+Table 3: Models used in our experiments, grouped by family and ordered by training data cutoff date.
 Model Family Model Variants Cutoff Date
-GPT-2 (Radford et al., 2019) GPT-2, GPT-2 XL 2017-12-01
-Pythia (Biderman et al., 2023) 14M, 160M, 2.8B 2020-03-01
-BLOOM (Scao et al., 2023) 560M, 7B1 2021-12-01
-LLaMA (Touvron et al., 2023) LLaMA-7B 2022-08-01
-OLMo (Groeneveld et al., 2024) 1B, 7B, 7B-Instruct 2023-03-01
-OLMo-2 (OLMo et al., 2024) 7B, 7B-Instruct 2023-12-01
-Table 3: Models used in our experiments, grouped by family and ordered by training data cutoff date.
-Inference.
-We use 1-shot in-context learning to provide models with a structured example of how to answer forecasting questions.
-For instruction-tuned models, we add an extra line of instruction to align with their training paradigm.
-To ensure fair comparison, baseline prompts are kept minimal while maintaining clarity.
-Confidence is estimated using logit-based normalized heuristics.
-Full prompts and hyperparameters are provided in Appendix C and  Appendix D.
+GPT-2 (Radford et al., 2019) GPT-2, GPT-2 XL 2017-12
+Pythia (Biderman et al., 2023) 14M, 160M, 2.8B 2020-03
+BLOOM (Scao et al., 2023) 560M, 7B1 2021-12
+LLaMA (Touvron et al., 2023) LLaMA-7B 2022-08
+OLMo (Groeneveld et al., 2024) 1B, 7B, 7B-Instruct 2023-03
+OLMo-2 (OLMo et al., 2024) 7B, 7B-Instruct 2023-12
+Qwen 2.5 (Yang et al., 2024) 1.5B-Instruct, 7B-Instruct 2023-12
+LLaMa 3.1 (Grattafiori et al., 2024) 8B 2023-12
+5.2 Inference
+Each model is evaluated independently using 1-shot in-context learning.
+For every question, we sample $n=10$ outputs from the model using temperature-based decoding.
+Each output includes a prediction $v_{i}$ and an associated confidence score $c_{i}$, derived from the normalized logits.
+Specifically, we compute the log-probability of each generated token using $\log p_{i,j}=\log\text{softmax}(z_{i,j})[x_{i,j}]$, where $z_{i,j}$ is the model’s pre-softmax logit vector at position $j$ for sample $i$, and $x_{i,j}$ is the generated token index.
+The confidence score $c_{i}$ is then defined as the average log-probability across all tokens: $c_{i}=\frac{1}{T}\sum_{j=1}^{T}\log p_{i,j}$, where $T$ is the length of the generated sequence.
+We observe low variance across the $n$ samples, with most models producing consistent predictions and confidence values.
 […]
-Model N Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$)
-GPT2 401 0.58 0.37 0.42
-GPT2-XL 401 0.67 0.32 0.45
-Pythia-14m 343 0.59 0.14 0.62
-Pythia-160m 343 0.61 0.26 0.53
-Pythia-2.8b 343 0.55 0.39 0.44
-Bloom-560m 263 0.49 0.41 0.45
-Bloom-7b1 263 0.64 0.33 0.32
-Llama-7b 226 0.57 0.34 0.52
-OLMo-1B 188 0.23 0.16 0.75
-OLMo-7B 188 0.21 0.17 0.82
-OLMo-7B-Instruct 188 0.65 0.14 0.38
-OLMo-2-7B 145 0.51 0.24 0.44
-OLMo-2-7B-Instruct 145 0.59 0.30 0.41
-Table 4: Performance of forecasting models on Boolean questions in FOReCAst.
-Reported metrics include the number of evaluated questions (N), accuracy, F1 score (both higher is better), and the Brier score (lower is better).
-5.2 Results, Findings, and Discussions
+Prompts are minimal and structured to ensure consistency across models.
+Each input includes a forecasting question and a single demonstration example.
+Instruction-tuned models receive an additional natural-language instruction to align with their training paradigm.
+Full prompt templates and decoding hyperparameters are provided in Appendix C and Appendix D.
+5.3 Results
 Our experiments on the FOReCAst dataset reveal that forecasting remains a challenging task for current LLMs, particularly in estimating and expressing uncertainty.
-While models often achieve reasonable accuracy in their point predictions, their confidence calibration—measured by Brier score and CRPS—shows substantial variability.
-This indicates that uncertainty estimation should be treated as a distinct evaluation axis, not inferred from accuracy alone.
-Throughout the following sections, we offer interpretations and possible explanations for the observed behaviors.
+While models often achieve reasonable accuracy in their point predictions, their confidence calibration, measured by Brier score and CRPS, shows substantial variability.
+This indicates that confidence estimation should be treated as a distinct evaluation axis, not inferred from accuracy alone.
+Table 4 reports results on a shared question set to ensure fair comparison across models, while Table 5 evaluates each model on all questions available up to its respective training cutoff as a view of each model’s forecasting potential.
+Table 4: Combined forecasting performance for cutoff date 2023-12, ensuring same questions for all models.
+CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
+Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
+GPT2 0.59 0.31 0.37 1.00 1.00 0.25 0.73 0.72
+GPT2-XL 0.71 0.30 0.42 1.00 1.00 0.04 0.69 0.64
+Pythia-14m 0.62 0.14 0.65 1.00 1.00 0.25 0.85 0.83
+Pythia-160m 0.67 0.31 0.51 1.00 1.00 0.02 0.67 0.65
+Pythia-2.8b 0.52 0.29 0.45 1.00 1.00 0.12 0.73 0.69
+Bloom-560m 0.52 0.38 0.41 1.00 1.00 0.03 0.67 0.65
+Bloom-7b1 0.68 0.36 0.30 1.00 1.00 0.11 0.78 0.75
+Llama-7b 0.54 0.27 0.55 0.96 0.94 0.06 0.62 0.59
+OLMo-1B 0.23 0.21 0.75 1.00 1.00 0.22 0.76 0.74
+OLMo-7B 0.22 0.23 0.83 1.00 1.00 0.22 0.82 0.81
+OLMo-7B-Inst 0.68 0.16 0.36 1.00 1.00 0.27 0.81 0.78
+OLMo-2-7B 0.51 0.24 0.44 1.00 1.00 0.02 0.65 0.62
+OLMo-2-7B-Inst 0.59 0.30 0.41 0.87 0.82 0.09 0.60 0.57
+Qwen-2.5-1.5B-Inst 0.67 0.20 0.32 1.00 1.00 0.14 0.74 0.72
+Qwen-2.5-7B-Inst 0.68 0.08 0.25 1.00 1.00 0.07 0.55 0.53
+Llama-3.1-8B 0.08 0.00 0.99 1.00 1.00 0.02 0.58 0.54
+Table 5: Forecasting performance across all tasks in FOReCAst, evaluated using questions available to each model based on its training cutoff date.
+As a result, the number of evaluated questions ($N$) may vary across models.
+Boolean question metrics include Accuracy, F1, and Brier score (higher is better for Accuracy/F1, lower is better for Brier).
+Quantity Estimation and Timeframe Prediction use lower-is-better metrics.
+APE: Absolute Percentage Error, MAE: Mean Absolute Error, CRPS: Continuous Ranked Probability Score, ADE: Absolute Days Error.
+Model Boolean Questions Quantity Estimation Timeframe Prediction
+N Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) N APE ($\downarrow$) MAE ($\downarrow$) CRPS ($\downarrow$) N ADE ($\downarrow$) CRPS ($\downarrow$)
+GPT2 401 0.58 0.37 0.42 81 0.23 0.87 0.86 26 0.99 0.99
+GPT2-XL 401 0.67 0.32 0.45 81 0.03 0.75 0.72 26 1.00 1.00
+Pythia-14m 343 0.59 0.14 0.62 77 0.17 0.88 0.85 25 1.00 1.00
+Pythia-160m 343 0.61 0.26 0.53 77 0.03 0.78 0.76 25 1.00 1.00
+Pythia-2.8b 343 0.55 0.39 0.44 77 0.07 0.81 0.79 25 0.97 0.96
+Bloom-560m 263 0.49 0.41 0.45 43 0.05 0.73 0.71 12 1.00 1.00
+Bloom-7b1 263 0.64 0.33 0.32 43 0.06 0.75 0.72 12 1.00 1.00
+Llama-7b 226 0.57 0.34 0.52 32 0.06 0.64 0.61 10 0.98 0.98
+OLMo-1B 188 0.23 0.16 0.75 23 0.21 0.79 0.76 6 1.00 1.00
+OLMo-7B 188 0.21 0.17 0.82 23 0.22 0.84 0.82 6 1.00 1.00
+OLMo-7B-Inst 188 0.65 0.14 0.38 23 0.24 0.80 0.77 6 1.00 1.00
+OLMo-2-7B 145 0.51 0.24 0.44 20 0.02 0.65 0.62 4 1.00 1.00
+OLMo-2-7B-Inst 145 0.59 0.30 0.41 20 0.09 0.60 0.57 4 0.87 0.82
+Qwen-1.5B-Inst 145 0.67 0.20 0.32 20 0.14 0.74 0.72 4 1.00 1.00
+Qwen-7B-Inst 145 0.68 0.08 0.25 20 0.07 0.55 0.53 4 1.00 1.00
+Llama-3.1-8B 145 0.08 0.00 0.99 20 0.02 0.58 0.54 4 1.00 1.00
 Boolean Questions
-Table 4 shows that while some models reach reasonable accuracy on binary classification tasks, their confidence calibration remains inconsistent.
+Table 4 and Table 5 show that while some models reach reasonable accuracy on binary classification tasks, their confidence calibration remains inconsistent.
 For instance, GPT2-XL achieves a relatively strong accuracy of 0.67 but has a higher Brier score of 0.45, whereas OLMo-7B-Instruct achieves a similar accuracy (0.65) with a lower Brier score of 0.38, indicating better alignment between predicted probabilities and actual outcomes.
 This discrepancy suggests that models may differ not only in their ability to identify the correct answer, but also in how they distribute confidence over binary options.
-Importantly, model performance does not appear to correlate straightforwardly with scale or training recency.
+Importantly, model performance, whether in terms of accuracy or confidence calibration, does not appear to correlate strongly with scale or training data recency.
 For example, LLaMA-7B, a more recent model, shows lower accuracy (0.57) and higher Brier score (0.52) than older models like GPT2-XL.
-This supports the observation that accuracy in binary forecasting tasks depends on a combination of world knowledge, probabilistic reasoning, and how well models can handle epistemic uncertainty, particularly when questions are unresolved or partially underspecified at prompt time.
-The gap between classification accuracy and calibration further underscores the importance of evaluating both, as good classification alone may mask unreliable probability estimates.
-Quantity Estimation Timeframe Prediction
-Model N APE ($\downarrow$) MAE ($\downarrow$) CRPS ($\downarrow$) Model N ADE ($\downarrow$) CRPS ($\downarrow$)
-GPT2 81 0.23 0.87 0.86 GPT2 26 0.99 0.99
-GPT2-XL 81 0.03 0.75 0.72 GPT2-XL 26 1.00 1.00
-Pythia-14m 77 0.17 0.88 0.85 Pythia-14m 25 1.00 1.00
-Pythia-160m 77 0.03 0.78 0.76 Pythia-160m 25 1.00 1.00
-Pythia-2.8b 77 0.07 0.81 0.79 Pythia-2.8b 25 0.97 0.96
-Bloom-560m 43 0.05 0.73 0.71 Bloom-560m 12 1.00 1.00
-Bloom-7b1 43 0.06 0.75 0.72 Bloom-7b1 12 1.00 1.00
-Llama-7b 32 0.06 0.64 0.61 Llama-7b 10 0.98 0.98
-OLMo-1B 23 0.21 0.79 0.76 OLMo-1B 6 1.00 1.00
-OLMo-7B 23 0.22 0.84 0.82 OLMo-7B 6 1.00 1.00
-OLMo-7B-Instruct 23 0.24 0.80 0.77 OLMo-7B-Instruct 6 1.00 1.00
-OLMo-2-7B 20 0.02 0.65 0.62 OLMo-2-7B 4 1.00 1.00
-OLMo-2-7B-Instruct 20 0.09 0.60 0.57 OLMo-2-7B-Instruct 4 0.87 0.82
-Table 5: Performance of forecasting models on Quantity Estimation (left) and Timeframe Prediction (right) tasks in FOReCAst.
-All metrics are lower-is-better.
-APE: Absolute Percentage Error, MAE: Mean Absolute Error, CRPS: Continuous Ranked Probability Score, ADE: Absolute Days Error.
+Moreover, LLaMA-3.1-8B fails entirely on this task, with an accuracy of 0.08 and a Brier score of 0.99.
+Manual inspection reveals that it often outputs malformed answers (e.g., long strings of zeros) rather than valid "Yes"/"No" responses, suggesting a failure to follow the prompt format.
 Timeframe Prediction
-Table 5 demonstrates that timeframe prediction—estimating when an event will occur—is especially difficult for current models.
+Table 4 and Table 5 demonstrate that timeframe prediction is especially difficult for current models.
 Across the board, both ADE and CRPS values remain close to 1.0, indicating little improvement over naive or uninformed baselines.
@@
 The generally poor performance suggests that temporal forecasting presents challenges beyond simple factual retrieval or classification.
-This difficulty likely reflects the inherent complexity of temporal prediction, which often requires models to reason over event dynamics, causality, and elapsed time—all aspects that are underrepresented in standard pretraining corpora.
+This difficulty likely reflects the inherent complexity of temporal prediction, which often requires models to reason over event dynamics, causality, and elapsed time, all of which are underrepresented in standard pretraining corpora, as well as the challenge posed by the broader space of possible answers, which extends far beyond binary classifications.
 In contrast to Boolean tasks, where resolution criteria are clearly defined and outcomes are binary, timeframe questions require mapping language input to a continuous temporal target, which demands a finer-grained understanding of how events evolve over time.
@@
 Quantity Estimation
-As shown in Table 5, quantity estimation—predicting a numerical outcome—exhibits a broader range of performance across models.
-Some models such as Pythia-160M and GPT2-XL achieve very low absolute percentage error (APE = 0.02 and 0.03 respectively), indicating strong point prediction capability.
-However, their CRPS values (0.76 and 0.72) suggest relatively poor calibration.
+As shown in Table 4 and Table 5, quantity estimation which predicts a numerical outcome exhibits a broader range of performance across models.
+Some models such as Pythia-160M and GPT2-XL achieve very low absolute percentage error (APE = 0.02 and 0.04 respectively), indicating strong point prediction capability.
+However, their CRPS values (0.65 and 0.72) suggest relatively poor calibration.
 In contrast, OLMo-2-7B-Instruct, with a higher APE of 0.08, attains the lowest CRPS (0.57), suggesting it better captures uncertainty even if its point estimates are less precise.
@@
 Since CRPS penalizes both distance from the true value and over/under-confidence, this metric provides a more comprehensive view of performance in numerical forecasting tasks.
-These results suggest that even relatively strong predictive models may require additional mechanisms—such as structured uncertainty modeling or calibration-sensitive objectives—to produce well-calibrated forecasts in quantity estimation settings.
+These results suggest that even relatively strong predictive models may require additional mechanisms such as structured uncertainty modeling to produce well-calibrated forecasts in quantity estimation settings.
+5.4 Findings
 Impact of Model Size on Forecasting Performance
 Model size alone does not reliably predict forecasting performance across tasks.
-Within the Pythia family, Pythia-2.8b occasionally improves over smaller variants in certain metrics—for example, it achieves lower CRPS than Pythia-14m on quantity estimation—but these gains are neither large nor consistent.
+As shown in Table 5, within the Pythia family, Pythia-2.8b occasionally improves over smaller variants in certain metrics.
+For example, it achieves lower CRPS than Pythia-14m on quantity estimation but these gains are neither large nor consistent.
 In some cases, smaller models like Pythia-160M outperform their larger counterparts in point prediction metrics such as APE.
@@
 These observations suggest that increases in parameter count do not automatically translate to improved forecasting performance.
-While larger models often benefit from greater representational capacity, the forecasting task—particularly when it involves expressing well-calibrated uncertainty—may place demands that are orthogonal to model size.
+While larger models often benefit from greater representational capacity, the forecasting task, particularly when it involves expressing well-calibrated uncertainty, may place demands that are orthogonal to model size.
 Specifically, forecasting tasks may require targeted reasoning over time, quantities, or epistemic uncertainty, which are not clearly emphasized during standard pretraining or scaled parameter growth.
 These findings indicate that effective forecasting may depend more on model behavior during decoding or task-specific adaptation, rather than scale alone.
-Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
-OLMo-7B 0.21 0.17 0.82 1.00 1.00 0.22 0.84 0.82
-OLMo-7B-Instruct 0.65 0.14 0.38 1.00 1.00 0.24 0.80 0.77
-OLMo-2-7B 0.51 0.24 0.44 1.00 1.00 0.02 0.65 0.62
-OLMo-2-7B-Instruct 0.59 0.30 0.41 0.87 0.82 0.09 0.60 0.57
 Table 6: Effect of instruction tuning on forecasting performance across boolean, timeframe, and quantity estimation questions.
 Metrics include accuracy, F1, and Brier score for binary questions; ADE and CRPS (T) for timeframe prediction; and APE, MAE, and CRPS (Q) for quantity estimation.
+Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
+OLMo-7B 0.21 0.17 0.82 1.00 1.00 0.22 0.84 0.82
+OLMo-7B-Inst 0.65 0.14 0.38 1.00 1.00 0.24 0.80 0.77
+OLMo-2-7B 0.51 0.24 0.44 1.00 1.00 0.02 0.65 0.62
+OLMo-2-7B-Inst 0.59 0.30 0.41 0.87 0.82 0.09 0.60 0.57
 Impact of Instruction Tuning on Forecasting Performance
@@
 0.82), indicating better confidence calibration.
-In timeframe prediction, the instruct-tuned OLMo-2-7B-Instruct improves uncertainty estimation, with an ADE of 0.87 and CRPS of 0.82, compared to 0.9981 and 0.9970 for the base OLMo-2-7B.
+In timeframe prediction, the instruct-tuned OLMo-2-7B-Instruct improves uncertainty estimation, with an ADE of 0.87 and CRPS of 0.82, compared to 0.9981 and 0.9970 (both rounded to 1.00) for the base OLMo-2-7B.
 For quantity estimation, instruct-tuned models have slightly higher APE but lower MAE and CRPS, suggesting better uncertainty calibration.
 While these results point to potential benefits of instruction tuning for probabilistic reasoning, the sample size remains limited due to the recency of instruction-tuned models, and the observed gains may not be statistically robust.
-Aggregation Method N Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$)
+Table 7: Different aggregation methods for extracting predictions and confidence from Llama-7B.
+Aggregation Method N Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$)
 Majority Vote 226 0.56 0.24 0.53
@@
 Bayesian Aggregation 226 0.58 0.35 0.55
-Table 7: Ablation study on different aggregation methods for extracting predictions and confidence from Llama-7B outputs.
 Impact of Aggregation Methods on Forecasting Performance
@@
 The challenge remains in developing techniques that optimize both accuracy and calibration simultaneously, emphasizing the importance of uncertainty-aware forecasting.
-In preliminary experiments, we also explored verbalized confidence, prompting the model to explicitly generate a probability (e.g., “I’m 80% confident…”).
+In preliminary experiments, we also explored verbalized confidence (Xiong et al., 2023), prompting the model to explicitly generate a probability (e.g., “I’m 80% confident…”).
 However, most models either failed to produce enough meaningful numeric estimates or generated inconsistent responses, particularly on harder questions.
 This highlights a limitation in current instruction-following capabilities when it comes to quantifying uncertainty directly.
-A promising direction for future work is to fine-tune models specifically to produce structured confidence outputs, possibly by supervising against calibrated targets or training with ranking-based objectives.
-Additionally, integrating task-specific priors (e.g., event resolution frequency, temporal proximity) into aggregation could further improve calibration.
-Hence, joint modeling of correctness and uncertainty remains a key open challenge for real-world deployment of forecasting systems.
+Therefore, future work includes fine-tuning for calibrated confidence and integrating task-specific priors.
+Joint modeling of accuracy and uncertainty remains a key challenge.
 Lack of Correlation Between Forecast Date and Model Performance
@@
 In many cases, model performance remains flat or even inconsistent across time, with no systematic degradation as the temporal distance increases.
-This observation suggests that forecasting is meaningfully different from retrieval-based tasks such as fact-checking or question answering.
+This observation suggests that forecasting is meaningfully different from retrieval-based tasks such as fact-checking or question answering (Guo et al., 2022).
 While those tasks often benefit from training data that includes the target facts or related content, forecasting involves reasoning under uncertainty about events that had not yet occurred during training.
@@
 These factors make the task more complex and reinforce the need for dedicated methods that go beyond static knowledge retrieval.
+These findings would not be observable under prior benchmarks that lack confidence metrics or diverse forecasting tasks.
 6 Related Work
@@
 7 Limitations
-FOReCAst is constructed entirely from Metaculus, a platform known for its high-quality, well-resolved forecasting questions and active user base.
-This makes it a strong and consistent foundation for developing the task of probabilistic forecasting with LLMs.
-However, it does not cover the full diversity of forecasting practices, such as different question formats, domain emphases, or forecasting styles seen in other platforms like Good Judgment Open or domain-specific expert settings.
-In this work, our primary focus is to define the task rigorously and establish a benchmark that is methodologically clean and scalable.
-Expanding to additional data sources and platforms is part of our future work and should help test generalization across forecasting cultures and environments.
-Our method for constructing gold confidence labels is based on community-aggregated forecasts, which offer a practical and scalable proxy for crowd belief and uncertainty.
-While this approach enables systematic evaluation across hundreds of questions, it may not fully capture the variability and richness of individual forecasters’ reasoning or levels of disagreement.
-Moreover, the aggregated predictions may sometimes be influenced by group dynamics or correlated sources of error.
-We acknowledge these trade-offs, and view our current approach as a reasonable starting point given the limited availability of high-quality, large-scale confidence annotations.
-Future work could explore incorporating richer uncertainty representations, such as forecast distributions, forecaster rationales, or disagreement metrics.
-Finally, our dataset is currently limited to English-language questions and forecasts, reflecting the language used by Metaculus and many existing NLP resources.
-While this is a reasonable initial scope for establishing the task, we recognize that forecasting behavior, uncertainty expression, and interpretation of probabilities may vary significantly across linguistic and cultural contexts.
-Extending FOReCAst to include multilingual questions and community forecasts is a promising direction for improving the inclusiveness and robustness of future benchmarks.
-We also note that some dimensions, such as temporal adaptation and real-time reasoning, remain out of scope here but are essential long-term goals for building forecasting-capable systems.
+FOReCAst is built entirely from Metaculus, a high-quality forecasting platform with clear resolution criteria and an active user base.
+This provides a strong foundation for defining and benchmarking probabilistic forecasting with LLMs.
+However, it does not reflect the full diversity of forecasting formats, domains, or styles seen in platforms like Good Judgment Open or domain-specific settings.
+Expanding to other platforms is an important next step to test generalization across forecasting cultures.
+Our scoring formulation aligns confidence with meaningful deviations from uncertainty and is tailored to Metaculus conventions.
+However, it can be adapted for use with other platforms or scoring protocols.
+The goal is to maintain an interpretable and consistent signal across all tasks.
+Gold confidence labels are derived from community-aggregated forecasts, offering a practical proxy for crowd belief and uncertainty.
+While this enables scalable evaluation, it may obscure forecaster disagreement and introduce biases from group dynamics or correlated errors.
+Future work could incorporate richer uncertainty signals, such as forecast distributions, forecaster rationales, or disagreement metrics.
+Finally, FOReCAst focuses on English-language questions.
+While this is consistent with Metaculus and existing NLP resources, probabilistic reasoning and uncertainty expression can differ across languages and cultures.
+Extending the benchmark to multilingual settings and modeling temporal adaptation and real-time reasoning are key directions for improving robustness and real-world applicability.
 8 Conclusion
@@
 Appendix B Dataset Statistics
+Table 8: Dataset statistics for FOReCAst, showing the distribution of questions across different forecasting types, with the overall total in the last column.
 Split Boolean Timeframe Quantity Total
@@
 Total 1758 139 349 2256
-Table 8: Dataset statistics for FOReCAst, showing the distribution of questions across different forecasting types, with the overall total in the last column.
 Table 9: Top 10 months and categories across the full dataset and each question type. Note that questions can belong to multiple categories, reflecting the diversity of domains represented.
@@
 Detailed model-specific performance metrics for Boolean Questions, Timeframe Prediction, and Quantity Estimation are presented in Table 10 to Table 14.
-Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
+These results highlight trends in forecasting accuracy and confidence calibration across models with different knowledge recency.
+Table 10: Combined forecasting performance for cutoff date 2017-12.
+CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
+Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
 GPT2 0.58 0.37 0.42 0.99 0.99 0.23 0.87 0.86
@@
 OLMo-7B 0.21 0.17 0.80 1.00 1.00 0.27 0.88 0.87
-OLMo-7B-Instruct 0.67 0.33 0.39 0.92 0.91 0.29 0.80 0.78
+OLMo-7B-Inst 0.67 0.33 0.39 0.92 0.91 0.29 0.80 0.78
 OLMo-2-7B 0.54 0.33 0.50 0.92 0.91 0.04 0.70 0.68
-OLMo-2-7B-Instruct 0.59 0.42 0.40 0.84 0.83 0.13 0.69 0.67
-Table 10: Combined forecasting performance for cutoff 2017-12-01.
+OLMo-2-7B-Inst 0.59 0.42 0.40 0.84 0.83 0.13 0.69 0.67
+Qwen-1.5B-Inst 0.67 0.32 0.32 1.00 1.00 0.07 0.73 0.71
+Qwen-7B-Inst 0.68 0.26 0.26 1.00 1.00 0.08 0.63 0.61
+Llama-3.1-8B 0.05 0.02 0.96 1.00 1.00 0.03 0.70 0.68
+Table 11: Combined forecasting performance for cutoff date 2020-03.
 CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
-Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
+Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
 GPT2 0.59 0.38 0.41 0.99 0.99 0.23 0.89 0.87
@@
 OLMo-7B 0.20 0.15 0.80 1.00 1.00 0.26 0.89 0.87
-OLMo-7B-Instruct 0.66 0.30 0.40 0.92 0.91 0.27 0.80 0.78
+OLMo-7B-Inst 0.66 0.30 0.40 0.92 0.91 0.27 0.80 0.78
 OLMo-2-7B 0.53 0.33 0.51 0.92 0.90 0.03 0.71 0.68
-OLMo-2-7B-Instruct 0.59 0.41 0.40 0.83 0.82 0.12 0.70 0.68
-Table 11: Combined forecasting performance for cutoff 2020-03-01.
+OLMo-2-7B-Inst 0.59 0.41 0.40 0.83 0.82 0.12 0.70 0.68
+Qwen-1.5B-Inst 0.66 0.32 0.33 1.00 1.00 0.07 0.75 0.73
+Qwen-7B-Inst 0.68 0.27 0.26 1.00 1.00 0.07 0.64 0.62
+Llama-3.1-8B 0.05 0.02 0.96 1.00 1.00 0.02 0.71 0.69
+Table 12: Combined forecasting performance for cutoff date 2021-12.
 CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
-Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
+Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
 GPT2 0.61 0.35 0.38 1.00 1.00 0.26 0.85 0.84
@@
 OLMo-7B 0.22 0.18 0.80 1.00 1.00 0.23 0.85 0.84
-OLMo-7B-Instruct 0.67 0.18 0.40 0.92 0.92 0.22 0.76 0.74
+OLMo-7B-Inst 0.67 0.18 0.40 0.92 0.92 0.22 0.76 0.74
 OLMo-2-7B 0.53 0.27 0.49 0.93 0.92 0.03 0.67 0.64
-OLMo-2-7B-Instruct 0.61 0.38 0.39 0.86 0.84 0.10 0.64 0.61
-Table 12: Combined forecasting performance for cutoff 2021-12-01.
+OLMo-2-7B-Inst 0.61 0.38 0.39 0.86 0.84 0.10 0.64 0.61
+Qwen-1.5B-Inst 0.66 0.24 0.33 1.00 1.00 0.10 0.75 0.72
+Qwen-7B-Inst 0.68 0.14 0.26 1.00 1.00 0.08 0.60 0.59
+Llama-3.1-8B 0.06 0.00 0.95 1.00 1.00 0.02 0.62 0.59
+Table 13: Combined forecasting performance for cutoff date 2022-08.
 CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
-Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
+Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
 GPT2 0.62 0.36 0.36 1.00 1.00 0.26 0.81 0.80
@@
 OLMo-7B 0.23 0.17 0.82 1.00 1.00 0.22 0.87 0.86
-OLMo-7B-Instruct 0.65 0.12 0.40 0.91 0.90 0.21 0.78 0.75
+OLMo-7B-Inst 0.65 0.12 0.40 0.91 0.90 0.21 0.78 0.75
 OLMo-2-7B 0.54 0.26 0.47 0.91 0.91 0.03 0.68 0.65
-OLMo-2-7B-Instruct 0.60 0.33 0.40 0.93 0.90 0.10 0.65 0.62
-Table 13: Combined forecasting performance for cutoff 2022-08-01.
+OLMo-2-7B-Inst 0.60 0.33 0.40 0.93 0.90 0.10 0.65 0.62
+Qwen-1.5B-Inst 0.68 0.25 0.31 1.00 1.00 0.11 0.76 0.74
+Qwen-7B-Inst 0.68 0.12 0.25 1.00 1.00 0.10 0.60 0.58
+Llama-3.1-8B 0.06 0.00 0.96 1.00 1.00 0.02 0.62 0.58
+Table 14: Combined forecasting performance for cutoff date 2023-03.
 CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
-Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
+Model Acc. ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
 GPT2 0.60 0.36 0.38 1.00 1.00 0.26 0.74 0.72
@@
 OLMo-7B 0.21 0.17 0.82 1.00 1.00 0.22 0.84 0.82
-OLMo-7B-Instruct 0.65 0.14 0.38 1.00 1.00 0.24 0.80 0.77
+OLMo-7B-Inst 0.65 0.14 0.38 1.00 1.00 0.24 0.80 0.77
 OLMo-2-7B 0.51 0.24 0.44 1.00 1.00 0.02 0.65 0.62
-OLMo-2-7B-Instruct 0.59 0.33 0.41 0.91 0.87 0.12 0.61 0.57
-Table 14: Combined forecasting performance for cutoff 2023-03-01.
-CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
-Model Accuracy ($\uparrow$) F1 ($\uparrow$) Brier ($\downarrow$) ADE ($\downarrow$) CRPS (T) ($\downarrow$) APE ($\downarrow$) MAE ($\downarrow$) CRPS (Q) ($\downarrow$)
-GPT2 0.59 0.31 0.37 1.00 1.00 0.25 0.73 0.72
-GPT2-XL 0.71 0.30 0.42 1.00 1.00 0.04 0.69 0.64
-Pythia-14m 0.62 0.14 0.65 1.00 1.00 0.25 0.85 0.83
-Pythia-160m 0.67 0.31 0.51 1.00 1.00 0.02 0.67 0.65
-Pythia-2.8b 0.52 0.29 0.45 1.00 1.00 0.12 0.73 0.69
-Bloom-560m 0.52 0.38 0.41 1.00 1.00 0.03 0.67 0.65
-Bloom-7b1 0.68 0.36 0.30 1.00 1.00 0.11 0.78 0.75
-Llama-7b 0.54 0.27 0.55 0.96 0.94 0.06 0.62 0.59
-OLMo-1B 0.23 0.21 0.75 1.00 1.00 0.22 0.76 0.74
-OLMo-7B 0.22 0.23 0.83 1.00 1.00 0.22 0.82 0.81
-OLMo-7B-Instruct 0.68 0.16 0.36 1.00 1.00 0.27 0.81 0.78
-OLMo-2-7B 0.51 0.24 0.44 1.00 1.00 0.02 0.65 0.62
-OLMo-2-7B-Instruct 0.59 0.30 0.41 0.87 0.82 0.09 0.60 0.57
-Table 15: Combined forecasting performance for cutoff 2023-12-01.
-CRPS (T) denotes the Continuous Ranked Probability Score for Timeframe Prediction, while CRPS (Q) denotes the Continuous Ranked Probability Score for Quantity Estimation.
-These results highlight trends in forecasting accuracy and confidence calibration across models with different knowledge recency.
+OLMo-2-7B-Inst 0.59 0.33 0.41 0.91 0.87 0.12 0.61 0.57
+Qwen-1.5B-Inst 0.65 0.20 0.34 1.00 1.00 0.13 0.72 0.69
+Qwen-7B-Inst 0.68 0.12 0.25 1.00 1.00 0.10 0.56 0.54
+Llama-3.1-8B 0.07 0.00 0.96 1.00 1.00 0.02 0.59 0.55